### v1

# Spatially–Structured Models of Viral Dynamics: A Scoping Review

Abstract
There is growing recognition in both the experimental and modelling literature of the importance of spatial structure to the dynamics of viral infections in tissues. Aided by the evolution of computing power and motivated by recent biological insights, there has been an explosion of new, spatially–explicit models for within–host viral dynamics in recent years. This development has only been accelerated in the wake of the COVID–19 pandemic. Spatially-structured models offer improved biological realism and can account for dynamics which cannot be well-described by conventional, mean-field approaches. However, despite their growing popularity, spatially–structured models of viral dynamics are underused in biological applications. One major obstacle to the wider application of such models is the huge variety in approaches taken, with little consensus as to which features should be included and how they should be implemented for a given biological context. Previous reviews of the field have focused on specific modelling frameworks or on models for particular viral species. Here, we instead apply a scoping review approach to the literature of spatially–structured viral dynamics models as a whole to provide an exhaustive update of the state of the field. Our analysis is structured along two axes, methodology and viral species, in order to examine the breadth of techniques used and the requirements of different biological applications. We then discuss the contributions of mathematical and computational modelling to our understanding of key spatially-structured aspects of viral dynamics, and suggest key themes for future model development to improve robustness and biological utility.
1 Introduction
There is a growing recognition in the virus dynamics literature of the role of spatial structure in influencing viral spread. Following the rapid evolution of computing power, and especially since the outbreak of the COVID–19 pandemic, there has been an explosion of new, spatially–explicit models for within–host viral dynamics. In line with the rapid advancement of sophistication and biological realism in modelling for viral infections, the incorporation of spatial structure enables a description of dynamics that cannot be well described by mean–field models, such as neighbour–based infection [123, 99, 173, 121, 215], or immune cell taxis [126, 111, 183, 78]. Spatially–structured models have also been shown to exhibit different dynamics to their mean–field counterparts under the same parameter values, suggesting that spatial structure may be necessary to improve estimates of biologically useful parameters [24, 85, 196].
However, despite their advantages, spatial models have been underused in an application context. Spatially–structured models of viral dynamics are routinely cited far less frequently and are much less commonly used in biological applications compared to mean–field compartmental models. While this may be in part due to the simplicity of the latter approach (including the difficulty in calibrating spatial models to experimental data), this is also likely due to the variety of approaches taken to modelling spatial structure in viral infections, and a lack of consensus on what features should be included and how they should be implemented. Mean–field models are more or less united in the common language of ODEs (possibly with some slight modifications to account for, for example, age structure or stochasticity), yet spatially–structured models span a huge variety of methods: from PDEs to compartmental approaches, to agent–based approaches, or combinations thereof. To add to the complexity, the actual nature of the spatial structure of infections by different viruses in different host tissues is also widely varied, depending on whether the infection is in vivo or in vitro, whether the host cells are motile or packed in tissue, how permissive the tissue is to virion diffusion, and so on. There is therefore a need for a wide–reaching, intersectional analysis of the full breadth of spatially–structured viral dynamics models to make sense of the breadth of techniques used and the requirements of different applications.
Earlier reviews of spatially–structured models of virus dynamics have generally focused on specific viruses or model methodologies. Graw and Perelson reviewed spatial effects in the dynamics of HIV and hepatitis viruses [85, 86]. Their work identified direct cell–to–cell spread of infection as a distinct motivator for the use of spatial models, and advocated for the use of agent–based models [86]. Bocharov et al., by contrast, focused on a PDE modelling approach specifically in HIV, and criticised a lack of models (as of 2012) which took into account tissue heterogeneity in HIV infection [39]. Bauer, Beauchemin and Perelson reviewed the use of agent–based models [24]. Their work highlighted the difference in dynamics seen in spatially–structured compared to mean–field models, and explored the challenges in calibrating and applying agent-based models [24]. They pointed out, that due to the motility of the target cells in HIV, this virus is somewhat of a special case, requiring a different approach to the treatment of spatial structure [24]. The most recent survey of the literature on modelling the spatial spread of viral infections was carried out by Gallagher et al. in 2018, in a review that mainly focused on influenza A infection [76]. As with Graw and Perelson, the authors explored mechanisms and implementation of the cell–to–cell infection route, and explained the key role this has in determining not only the spatial structure but also the global dynamics of infection [76]. In contrast to Graw and Perelson, the authors advocated against the use of agent–based modelling at least for the immune response, preferring instead a compartmental approach [76].
While these early reviews provided an important overview and insight into approaches to modelling the spatial structure of viral infection in varying domains, there has so far yet to be a review spanning the full scope of modelling methods and viral species. Moreover, there is a need to take stock of the rich body of work that has emerged in recent years, especially following the emergence of SARS–CoV–2. Here, we employ a scoping approach to the literature on spatially–structured models of viral dynamics in order to systematically and exhaustively review the full range of approaches used.
The remainder of this paper is structured as follows. In Section 2 we present the methodology used for our scoping review, including inclusion / exclusion criteria. We then present an overview of the key publications in Section 3, before moving on to our detailed review of the field. In Sections 4 and 5 we present the state of the field by methodology and virus species respectively, before discussing innately spatially dynamic models in Section 6. Finally, in Section 7, we present emerging trends in the future developments of spatial modeling of viral dynamics, before closing the paper in Section 8.
2 Methods
2.1 Search string
We conducted a search of the literature on September 2nd, 2024, using the search query
(“virus” OR “viral”) AND (“dynamicOR “kinetic”) … x                            AND (“spatiOR “space”)… x                            AND (“mathOR “compOR “sim”)… x                            AND “model
where AND and OR are the usual boolean operators and is the wildcard token. We queried Scopus and Web of Science using this search string against article titles, abstracts and keywords. We also queried the search string against Scopus’s pre–print database (this functionality was not available from Web of Science).
2.2 Inclusion and exclusion criteria
After merging results from both databases and screening for duplicates, we analysed the combined pool of studies and pruned for relevance. Our criterion for inclusion was studies that introduce or apply a spatially–explicit mathematical or computational model of within–host viral dynamics. We excluded articles that met one or more of the following criteria:
  • Review articles. We omitted review articles from our search such as those mentioned in the previous section [86, 85, 76, 39, 24], preferring instead to limit our focus to original modelling applications.
  • Studies focused purely on within–cell dynamics. Some studies modelled only the within–cell life cycle of viruses, especially hepatitis B and C [79, 150, 114]. Since our focus is on models of viral infection spread between host cells, we excluded these studies from our analysis.
  • Studies without an explicit viral or infected population. This includes some computational models which described only the immune system, sometimes with a general “antigen” population [51, 88, 162].
  • Studies with a mainly analytical focus. Our search returned a substantial body of studies focused on the mathematical properties of spatial models of viral dynamics such as existence and smoothness of solutions, stability properties and conditions for Turing patterns (in fact, our database search found more studies that met this description than studies that were retained for analysis). While these works offer a substantial contribution to the collective literature on mathematical modelling of viral infections and are worthy of review in their own right, our work here is focused primarily on the application of such models and as such, we excluded these studies from our analysis. Note that while the determination of a study as “analytically–focused” as opposed to “application–focused” is difficult to articulate objectively, we found that the distinction between works with a mathematical readership in mind and those with a more interdisciplinary audience was usually very clearly determined. In the few instances where this distinction was unclear, we opted for including the article.
We conducted three iterations of screening of the studies found by our search, applying the inclusion criterion to first the titles, then the abstracts, then the full text of the remaining studies, discarding those which clearly did not meet the inclusion criterion. We only applied the exclusion criteria at the full text level. Of the 3492 unique studies returned by our search, we initially found 92 that were relevant to this review. The largest subset of publications which were removed on the basis of the exclusion criteria above were articles which were deemed to have an analytical focus (we found 111).
After having applied our selection criteria, we then screened the references of each of the accepted articles to find any articles missed by our database search. This process found a further 84 relevant studies for a total of 176. We outline the general workflow of the literature search process in Figure 1. We conducted our search in English and only considered articles written in English. We included pre–prints in our search.
Figure 1: Workflow for the scoping review. Duplicates automatically detected in JabRef. Duplicates manually detected.
2.3 Categorisation
In order to analyse high–level trends in the literature, we collected for each accepted study the viral species modelled (where one was specified) and the model framework used. We considered continuum models, patch models, cellular automata, and multicellular models in this review, which we define as follows. Models in which the spatial domain and all model components are continuous in space (that is, they are represented by densities, rather than discrete agents) are classified as continuum models. These are overwhelmingly PDE models, especially reaction–diffusion systems. Of the models which include spatially–discrete regions, we define patch models as those which represent cells as a density. Such models typically comprise a lattice of connected spatial regions, each of which contains a (possibly stochastically–simulated) ODE system. Models which included only a few spatial regions which were hard–coded to represent specific organs or tissue regions were not included in our search. The remaining models to be classified comprise spatially–discrete regions and a discrete number of cells and can collectively be considered agent–based models. We subdivide these models into cellular automata and multicellular models based on whether the model considers a single spatial scale or multiple, respectively. For instance, a model in which the cells are taken to be discrete spatial regions and where the extracellular viral load is tracked as the quantity of viral density at each cell would be considered a cellular automaton, since the cell–scale is the only spatial scale considered by the model. By contrast, if extracellular viral load were represented as an overlying field governed by a PDE, for example, the model would be considered multicellular. Note that we consider models with a global virus reservoir which has equal access to all spatial regions as having a secondary spatial scale and therefore multicellular. We summarise this classification in Figure 2. We assign to each study a single modelling framework: for those which make use of multiple models of different frameworks, we categorise the article by the model with the highest level of spatial resolution used (according to our classification, which corresponds to the furthest to the right on the tree in Figure 2). For each study, we also recorded its number of citations according to Google Scholar as of September 2nd, 2024.
Figure 2: Classification for modelling frameworks discussed in this review. Note, the boundaries of the methodology classifications used here are maintained through out the reminder of this paper.
3 An overview of the use of spatially–structured models in viral dynamics
Our search procedure returned 176 publications, spanning 60 years of research across a diverse group of virus species and a wide array of methodologies. The final annotated list of publications retained for analysis can be found on FigShare [^1] https://doi.org/10.26188/27085984.v1. In Figure 3 we present a broad overview of the distribution of the studies found by our search over time, and by virus modelled and method used. Figures 3(a)–(b) shows the averaged number of publications per year over time, computed by taking the average number of publications per year across the five–year interval centred on a given year. For 2022 and 2023 we compute the average over a shorter interval (from two years prior, up to 2023). We do not show data from 2024. In Figure 3(a) we stratify these studies by virus species modelled and in Figure 3(b) we stratify by methodology. In Figure 3(c), we show the overall split of methodologies used and for which virus species.
To our knowledge, the earliest spatial description of viral spread in tissue with mathematical models was by Koch in 1964, who used the diffusion (heat) equation to analyse the growth of viral plaques in vitro [115]. However, the first “model” in a more recognisable sense, and in some ways the birth of the field of spatially–structured modelling for viral dynamics, was proposed some decades later by Pandey in 1989 [157]. Pandey’s model used a cellular automaton approach for the dynamics of HIV, and most of the spatial viral dynamics models in the early 1990s followed in this vein, as can be seen in Figures 3(a)–(b) [157]. Figures 3(a)–(b) also show significant growth and diversification of the literature on spatial spread of viral infections in the first half of the 2000s, with the first spatially–structured models specifically for influenza, hepatitis B and C (HBV and HCV), oncolytic viruses and others, along with the first patch and multicellular models all published around a similar time. Since 2020 and the outbreak of the COVID–19 pandemic, there has been a spike in the development of spatial models for viral infections, with the average yearly number of publications roughly doubling from less than six in the mid–’10s to approximately 13 in 2023. Curiously, this cannot be entirely attributed to models for SARS–CoV–2, with models for the virus accounting for less than a third of recent publications (19 of 62 publications since 2020).
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
Figure 3: Spatially–structured viral dynamics models over time (a) by virus species and (b) by methodology. (c) Overall split of models by methodology.
As shown in Figure 3(c), continuum models formed the largest proportion of the studies found by our search with 56, although between cellular automaton (54) and multicellular (52) approaches, we observed that agent–based models in general were a more popular approach (106 publications combined). Early agent–based models were almost exclusively cellular automata, although this approach fell out of favour in the 2010s, where multicellular approaches rose to prominence, as can be seen in Figure 3(b). Patch models were the least popular approach with only 14 publications since 1985. These were mostly models for HIV and herpes simplex virus 2 (HSV–2) with six publications each, including the entire spatial modelling literature specifically focused on HSV–2. The other modelling approaches were also highly specific to particular viral species with the largest proportion of continuum models focused on oncolytic viruses and the largest share of cellular automaton models developed for HIV. Multicellular models, by contrast, were most commonly proposed for general viral infections and not a particular viral species.
We sought to identify which of the publications found had been the most impactful. To account for the time since publication, we quantified impact by considering both a study’s overall number of citations, and the average number of citations it had received per year since publication, and imposed thresholds on both quantities. These constraints ensure that a publication has both reached a specified degree of influence overall, but also that it has been influential relative to the number of years since its publication. In Figure 4, we plot each publication found in our search against its overall number of publications and its average number of publications per year. Publications are colour–coded by virus and markers indicate methodology used. We opted for thresholds on the overall and per–year citation counts of 50 and 6, respectively, and found 25 studies that passed both thresholds. We show a timeline of these seminal publications in Figure 5. In the proceeding analysis, we make particular mention of these high–impact publications. In this review, articles that were included in our search are referenced alongside the authors’ names, and high–impact publications are indicated in bold.
Figure 4: Selection of seminal papers. We show each accepted study for our scoping review by total number of citations and average citations per year since publication. Cutoffs for inclusion in our analysis are shown.
Figure 5: Seminal spatially–structured viral dynamics models over time. Criteria for inclusion are more than 50 total citations and more than 6 citations per year on average since publication. Pre–print. Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Wu et al. [212], Wein et al. [205], Funk et al. [75], Beauchemin et al. [27], Friedman et al. [73], Xiao et al. [213], Mok et al. [146], Heilmann et al. (2010) [93], Mitchell et al. [145], Wodarz et al. [210], Yakimovich et al. [215], Heilmann et al. (2012) [94], Schiffer et al. [177], Bocharov et al. [38], Malinzi et al. [132], Friedman and Lai [72], Eriksen et al. [66], Imle et al. [99], Alzahrani et al. [9], Getz et al. [78], Quirouette et al. [166], Simmons et al. [190], and Sego et al. [183].
4 Spatially–structured models by methodology
We begin our analysis by discussing the modelling methodologies used across the literature on the spread of viral infections in space. The choice of methodology used in a given model has significance beyond just the modeller’s preference: depending on the length scale of interest, or the specific dynamical features to be explored, certain modelling frameworks may be more, or less suitable. Moreover, there is an inherent trade–off between simpler models — which are more amenable to analysis and more readily calibrated to experimental data — and larger, more detailed models, which offer greater biological precision and realism. As such, a model’s structure and design is essential to understanding the dynamics that model predicts, and the specificity and robustness of inferences made from such a model. In the following section, we discuss the methodological approaches used within the four modelling categories described in Section 2, that is, continuum models, patch models, cellular automata and multicellular models.
4.1 Continuum models
Figure 6: Continuum models summary. Seminal publications are as follows: Wu et al. [212], Wein et al. [205], Friedman et al. [73], Mok et al. [146], Bocharov et al. [38], Malinzi et al. [132], Friedman and Lai [72], Alzahrani et al. [9], and Quirouette et al. [166].
Continuum models represent the spatial domain and all model components as continuous in space. For the models we found in our search, this meant systems of PDEs, possibly coupled to ODEs, governing the dynamics of density distributions of cells, viruses, cytokines, and so on. The methodological simplicity of this approach also makes continuum models conducive to analytical approaches not possible under other frameworks, such as travelling wave and stability analysis. In representing all model components — and cells in particular — as density distributions in space, such models implicitly assume a macroscopic length scale relative to the size of the cell, yet the continuous spatial domain still offers a fine degree of spatial resolution. These properties make continuum models a popular choice for modelling the diffusion and possibly advection of virus and cytokine densities, for example, and the mechanical properties of tissues at the whole–tissue scale. It is perhaps for this reason that continuum models are a particularly popular modelling choice for oncolytic viruses (21 of the 56 continuum models found), where it is of interest to track the the spread of infection through cancer cells at tumour–scale, and also to take into account the mechanical properties of the growing tumour.
The simplest continuum models use a single PDE to track viral spread in space. We mentioned in the previous section Koch’s analysis of plaque spread using properties of the diffusion equation [115]; similar approaches have since been employed by others. Rodrigo et al. also used properties of the diffusion equation, in addition to a stochastic model for multiple infection foci, to study the growth of plaques in viral infections of plants [170]. Baabdulla et al. used a homogenised Fisher–KPP equation to analyse the spread of influenza A viral load across a tissue culture comprising a structured mix of infection–prone and infection–resistant cells, and found that a good fit was obtained only when the structure of the tissue was well–represented in the homogenisation process [18]. Bocharov et al. also employed a Fisher–KPP model for the spread of viral density in permissive tissues, with the addition of a general time–delayed, immune–mediated decay term [38]. Even with this simple model, the authors demonstrated complex travelling wave dynamics whose speed and structure depend on the size of the initial viral inoculum [38]. Moreover, the authors demonstrated the possibility of spatially–homogeneous or oscillatory dynamics depending on the properties of the viral diffusion and immune response [38].
Many continuum models are very closely related to well–established ODE systems. Such models bear the same target cell limited structure as many ODE systems but in spatially–explicit conditions, mediated by the spatial spread of one or more model components. Often, this is simply encoded via viral diffusion: Frank [71] used this approach to study the spread of replication–competent and defective–interfering viruses; others have used such models to track the spreading speed of infections (Ait Mahiout et al. [1], Amor and Fort [11], de Rioja et al. [59]), possibly in the context of competing viral strains (Ait Mahiout et al. [2], Tokarev et al. [201] and Holder et al. [96]). In some contexts, cell density is also assumed to diffuse in space. Alzahrani made this assumption in a model of SARS–CoV–2 infection where both healthy and infected cells were assumed to be motile [8], although diffusion of cell density is more commonly found in models of oncolytic viruses, where cellular diffusion accounts for the expansion of the tumour in space (Alzahrani et al. [9], Friedman and Lai [72], Malinzi et al. [132], for example).
Model dimensionality is also an important consideration for the dynamical and analytical properties of continuum models. Models with a single spatial dimension have been popular since, despite their simplicity, a single spatial variable is often sufficient to describe a key axis of spatial heterogeneity. Quirouette et al. [166], Vimalajeewa et al. [203] and Ait Mahiout et al. [1] have all employed PDE models of respiratory viral infections with a single spatial variable representing depth along the respiratory tract; Holder et al. [96] and Haseltine et al. [92] have used radial spatial variables to describe the growth of radially–symmetric viral plaques. In models of oncolytic viruses, the three–dimensional tissue environment is frequently collapsed down to a single spatial axis, either representing a travelling front of cancer cells along one spatial dimension (Malinzi et al. [131, 133]), or distance from the centre of a radially–symmetric tumour (Wu et al. [212], Wein et al. [205], Friedman et al. [73], for example). Dunia and Bonnecaze also used this radial symmetry approximation but to represent generalised viral infection in spherical organs [64]. One–dimensional models have the additional benefit of being more amenable to analytical approaches, including the calculation of wave speeds for the infection front (Bocharov et al. [38], de Rioja et al. [60], for example). Amor and Fort [11, 12], Fort and Méndez [70] and Bessonov et al. [34] have also both shown that delays associated with the within–cell viral life cycle strongly impact the rate of speed of the infection wave. Beyond a single spatial dimension, others have also modelled the spread of infections using continuum methods in two dimensions (Graziano et al. [87], Alzahrani et al. [9], Reisch and Langemann [169], Liang et al. [128], for example) and, less commonly, three dimensions (Friedman and Lai [72], Bailey et al. [20]). The advantage of the higher–dimensional (and higher–complexity) approach is the ability to capture anisotropic infection distributions and more detailed spatial heterogeneity than is possible in a single spatial dimension. Liang et al. employed a reaction–diffusion system with stochastic effects at low density to study the role of defective interfering particles in viral infection spread, and demonstrated the formation of symmetry–breaking “patchy” infection plaques [128]. Mbopda et al. showed varying spatial patterning as a function of the viral diffusion parameter in a PDE model of hepatitis B dynamics [140].
In addition to diffusion, continuum models have also regularly been used to describe anisotropic transport of both viruses and cells. Anekal et al. analysed the dispersal of virions under the influence of fluid flow in two spatial dimensions, and showed the formation of a variety of spatial patterns depending on the adsorptive properties of the virions and the characteristics of the fluid [13]. Directed transport is also a key feature in models of oncolytic viral dynamics due to the moving and growing environment. Timalsina et al., for example, included advection of the components of their model with the spread of tumour density [199]. Another key aspect of oncolytic viral dynamics is the directed migration of both cells and free virions towards regions of higher extracellular matrix density, known as haptotaxis. Alzahrani et al. [9, 10] and Alsisi et al. [6, 7, 5] have described in depth the role of extracellular matrix density and haptotaxis in determining the spreading pattering of tumours under oncolytic virotherapy. In brief, these works collectively show that haptotaxis can function to break the symmetry of the spreading tumour, and moreover that this hampers and directs the spread of free virions [9, 10, 6, 7, 5].
There are some limitations to the continuum modelling approach. As Graw and Perelson pointed out in their review, continuum models — both PDE and ODE systems — fall under the umbrella of mean field approximations, in the sense that interaction between model species in such a model framework requires co–location, and cannot take into account interaction between spatially–adjacent quantities, such as direct viral spread between neighbouring cells [85]. It is also difficult to draw sharp boundaries to infected regions using a continuum approach, although these are clearly observed biologically: the usual assumption of linear diffusion for, say, the virus population results in the instant transport of a low level of viral density and therefore infected cell density everywhere in the spatial domain. Explicit moving tumour boundaries have been applied in models of oncolytic viral infections by Alzahrani et al. [9, 10], Alsisi et al. [6, 7, 5], Timalsina et al. [199] and Kim et al. [113] as a means of overcoming this limitation. However, even in these cases, the moving external boundary does not account for the boundaries of infected regions which are internal to the tumour. The moving boundary approach moreover does not allow for the formation of multiple, disparate infected regions, which are observed experimentally [74, 134, 41]. Another shortfall of the continuum approach is its limitation in describing rare stochastic events, such as extinctions or mutations, which occur at low densities. Baabdulla and Hillen, for example, observed oscillations in tumour load in a PDE model for oncolytic virotherapy that occasionally reached very low levels, but, due to the continuum framework of the model, never resulted in extinction [17]. Some authors have implemented stochastic elements to models otherwise described by PDEs in an aim to include account for low–density phenomena. We have already mentioned the approach taken by Liang et al. in stochastically simulating low–density viral and DIP dynamics [128]; Li et al. also used a semi-stochastic scheme for a PDE model to account for viral mutation in HCV infection [127]. Bailey et al. took a different approach and used a mainly probabilistic model to quantify the efficacy of randomly–distributed, radially–growing infection foci in oncolytic virotherapy [20]. While this approach sacrifices much of the mechanistic properties of other continuum models (the size of the infected population per plaque is assumed to grow at a constant rate, for example), the trade–off is a quantification of the inherent variability in efficacy of oncolytic virotherapy [20].
4.2 Patch models
Figure 7: Patch models summary. Seminal publications are as follows: Funk et al. [75], and Schiffer et al. [177].
Patch models divide the spatial domain into discrete, interconnected regions (or “patches”), within each of which conditions are assumed to be well–mixed. Since cell populations under the patch framework are represented as densities (as opposed to discrete counts, which we defined as the distinction between patch and cellular automaton models), this construction relies on each patch containing sufficiently many cells that the density assumption remains valid. As such, patch models, like continuum models, are best suited for macro–scale modelling. In their simplest and most common form, patch models are defined as a network of regions, each governed by a set of ODEs, where connectivity between the regions is implemented through the flow of one or more model components, usually virus density. Assuming the model governing each patch is identical aside from the patch connectivity, such models can be thought of as closely related to both PDE models — in a sense, this setup is a reaction–diffusion PDE system with spatial averaging over the patch regions — and also mean–field ODE models: the overall dynamics of a patch system are governed by the within–patch model with the addition of a spatial delay as the infection spreads between patches. The earliest patch models were proposed by Orive et al. [154] and Funk et al. [75], both of whom derived key results about the role of infection spread in patch networks. Orive and colleagues showed that increased migration between patches reduced the chance of the infection taking hold [154]. Funk and coworkers found that extent of coupling between compartments affected the global rate of infection, as well as the long–time dynamics of the system [75]. The authors demonstrated conditions under which the ability of the system to form an equilibrium depended on the extent of viral flow between compartments [75].
The patch approach can be considered an extension of compartmental ODE models. For example, some authors have considered distinct ODE systems for the upper and lower regions of the respiratory tract, where the two compartments are coupled by the flow of virions [112, 81]; others considered more complex networks of interconnected organ compartments [204] (note that we do not consider such models explicitly spatial and thus did not include these in our analysis). Patch models extend this approach to an arbitrary network or lattice of compartments. Among the patch models found by our search, a popular approach was to employ a regular structure to the patches, with either rectangular (Funk et al. [75], Byrne et al. [43], Zhang et al. [223]) or hexagonal packing of the patches (Schiffer et al. [179, 180, 177, 178, 176, 61]). The patch approach has the property that its spatial structure is only defined implicitly via the strength and network of the connectivity between the patches. As such, depending on the structure of the patch connectivity, such models can reflect complex biological topologies which are not easily described by continuum models, for example. Nakaoka et al. proposed a patch model for the spread of HIV infection between lymph nodes, explicitly representing the complex interconnectivity of the lymphoid tissue network, while treating the dynamics within each lymph node as well–mixed [151]. Cardozo et al. [48, 47] and Jagarapu et al. [105], by contrast, applied a patch model at the within–lymph node scale. Their model considered distinct lymphoid follicles, interconnected via a free blood and lymph compartment, where the spherical lymphoid follicles were in turn subdivided into radial shell patches [105, 48, 47].
The within–patch dynamics of patch models need not be governed by deterministic systems. A model put forward by Schiffer et al. and used in a number of applications for HSV–2 infection considered a hexagonally–packed lattice of patches governed by stochastically–simulated ODE systems [179, 180, 177, 178, 176, 61]. The authors employ stochasticity due to the random nature of the initiation of viral shedding and clearance [176]; as discussed above, these low–density phenomena are not well–captured by continuum approaches. Another stochastic patch system was proposed by Zhang et al. for the dynamics of HIV infection [223]. In their model, which borders on a cellular automaton approach, infection and death of cells are modelled as probabilistic events which occur between virions and cells in the same lattice site, while diffusion is rendered as a deterministic, continuous flow between adjacent compartments [223].
The patch modelling approach is also subject to shortcomings, and is by far the least used methodology for spatial viral dynamics modelling among the frameworks we considered. ODE patch models suffer from the same drawback as continuum models in that the continuous flow of virions between compartments prevents the determination of distinct boundaries to infected regions. In fact, the boundaries of infected regions are even less clearly defined for patch models since these models only possess spatial resolution at the patch scale. On the other hand, patch models do not have the benefit of being especially conducive to mathematical analysis. Some theory for within–host patch modelling was developed by Orive et al. [154] and Funk et al. [75], although only for two–patch populations or approximate models. Moreover, with the exception of network–structured models, such as that of Nakaoka et al. [151], the regions of patch models do not have any specific meaning. These patches instead represent somewhat arbitrary mesoscopic regions of cells which are large compared to the size of the cell (such that a density description is valid), yet small compared to the overall size of the tissue (in order to impose spatial structure).
4.3 Agent–based modelling: general approaches
The class of models characterised by the presence of discrete spatial regions and a discrete cell population (these spatial regions may or may not be the cells themselves) can broadly be described as agent–based. We consider two sub–types of agent–based models in this analysis: cellular automata, where there is only one spatial scale present, and multicellular models, which include more than one spatial scale. These two sub–types have unique nuances which we articulate in the following sections, but there are substantial overlaps between the two approaches which we discuss here.
Agent–based models which treat individual cells as spatially–discrete regions permit the possibility of integrating the within–cell viral dynamics of each cell. This is generally incorporated using a patch–like approach, where each cell is assigned an ODE system governing its spatially–homogeneous internal dynamics (Akpinar et al. [3], Sego et al. [183], Goyal and Murray [82], Bouchnita et al. [40], for example). Note that incorporating within–cell dynamics in this way does not explicitly add an additional spatial scale, and thus can be implemented in a cellular automaton approach as we have defined it (Akpinar et al. [3] does so), however, within–cell dynamics are more commonly found in more complex, explicitly multiscale multicellular models. There are a range of benefits from including within–cell viral dynamics. For one, explicitly representing the within–cell viral life cycle offers a more mechanistic account of the well–established lag observed between the time at which a cell is infected and the time at which a cell begins releasing virions (the eclipse phase), usually accounted for by a blanket latently–infected phase [183, 192, 218, 95]. This added detail moreover allows for a more detailed description of virus uptake and release by target cells (Sego et al. [183], Getz et al. [78], Rowlatt et al. [172]). Virus uptake is otherwise usually accounted for only under a general viral “decay” term [139, 23, 37, 192, 218]. Furthermore, in accounting for components of the viral life cycle in more detail, models can be made to more closely represent experimental systems. In hepatitis virus dynamics, for example, multiple sources of experimental data track quantities of intracellular components such as intracellular viral RNA or covalently closed circular DNA (cccDNA) and models that account for such quantities are therefore better equipped to model the experimental findings [86, 82].
A popular approach in building agent–based models for viral infections has been in the development of generalised, often highly complex computational platforms which are modular in structure, such that a core modelling framework can readily be extended to incorporate the specifics of a particular application. Such platforms are sometimes self–described as “simulators” instead of models. One of the earlier instances of such an approach was adapted from an immune simulator proposed by Caleda and Seiden [51] (in its original statement, the model did not include an explicit viral population and was therefore not included in our analysis). This was model was later developed by the authors and others — including Bernaschi and Castiglione — into the “ImmSim” platform [49, 35, 116, 33, 50, 21], and an offshoot, “ParImm” [32]. The central model of these systems describes an exceptionally complex virtual immune system powered by a so–called “bit–string” methodology, where each all modelled molecules, virions, and cell binding sites are represented as binary strings, and where interaction between these agents is governed by similarity between strings [33, 32] (this bit–string approach is rarely used among the spatial viral dynamics literature, although Zhang et al. [223] and Fachada et al. [67] used similar ideas). The underlying immune system model of ImmSim and its derivatives has been adapted to be applied to generalised viral infections [35, 116] or specific viral species such as HIV [33, 50, 21] or Epstein–Barr virus (EBV) [49]. In addition to ImmSim, Meier–Schellersheim and Mack proposed an alternative cellular automaton immune simulator, “SIMMUNE,” although, to our knowledge, this has not been applied to modelling viral infections since the original statement of the model [141]. Yakimovich et al. developed another modelling platform called “Infectio”, which coupled a discrete cell grid to a particle strength exchange model for viral flow [214, 215]. Other authors have built agent–based simulation models catered to more specific scenarios. Shapiro et al. developed a computational model for EBV infection called “PathSim” which explicitly incorporated the structure of Waldeyer’s ring [187, 63], and more recently, Moses et al. proposed “SIMCov”, a model for SARS–CoV–2 infection in a virtual human lung [149].
In recent years, large simulator models have again become popular, although recent approaches are more commonly multicellular in structure and account for multiple spatial scales. Perhaps the most substantial among these was proposed by Getz et al., which introduced an enormously detailed “tissue simulator” for SARS–CoV–2 infection [78, 101]. This model incorporated a detailed description of virus–host interactions, as well as a complex immune system, within–cell dynamics, and even tissue mechanics over time [78, 101]. Moreover, the model was developed with the explicit intention of modularity, such that further sub–models might be later incorporated into its architecture [78]. An alternative multicellular simulator of acute respiratory infection, coined the “Cellular Immunity Agent–Based Model (CIABM)” was suggested by An et al., which the authors used to demonstrate heterogeneous viral dynamics between different hosts [29, 56]. Rowlatt et al. proposed a similarly detailed model of respiratory infection, which they used to isolate the role of specific immune components in SARS–CoV–2 infection, especially macrophages [172]. Recent modular platforms have increasingly been integrated with existing software packages for tissue dynamics. The Getz et al. model [78, 101] was developed in PhysiCell [80], for instance, while a recent series of models by Glazier and colleagues have been developed mainly with CompuCell3D [182, 183, 184, 14, 68]. Among these, one model by Sego et al. [183] explicitly demonstrated cross–platform compatibility by being implemented in CompuCell3D [197], and also being implemented in Chaste [143, 163], and Morpheus [194] for later comparison work.
In all, these highly complex computational models serve as simulation “engines”, such that modellers or other end–users can simulate and visualise infection dynamics with a high degree of biological detail, with the possibility of building in further layers of model detail specific to a given application. The drawback to this approach is that the high degree of complexity significantly increases the difficulty of analysis, both of the dynamics of the model, but also of parameter identifiability. This is exacerbated by the enormous computational costs associated with these models, which can include millions of individual agents tracked over complex spatial domains. As such, for some of these models it is difficult to study the interaction of different components to any degree of detail, and virtually impossible to conduct statistically robust parameter estimations for such models (if sufficient data were even to exist). Parameter values for such models can therefore only be extracted from other models which may have varying relevance to the model to which they are applied. Clearly, there is a distinct application in mind for such models compared to, say, lightweight mean–field models for viral dynamics. While immune and tissue simulators are useful for hypothesis generation and simulation, for these to be useful for precision exercises such as parameter estimation, further advances in computational and statistical inference techniques are needed.
4.4 Agent–based modelling: cellular automata
Figure 8: Cellular automata summary. Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Beauchemin et al. [27], Xiao et al. [213], Heilmann et al. (2010) [93], Wodarz et al. [210], and Heilmann et al. (2012) [94].
Cellular automata are a class of discrete computational models. Like patch models, cellular automata involve discretising the spatial domain into a generally regular grid of explicit regions, however, the point of departure is that in cellular automata, the cell population is also discrete. Since this modelling approach involves tracking individual cells, cellular automata are commonly used for meso– or microscopic spatial scales, although cellular automata with vast cell populations are known (Moses et al. [149] and Shapiro et al. [187, 63], for example). Cellular automata are governed by distinct, often stochastic and adjacency–based rules, and as such generate dynamics which are notably distinct from those of continuum models. Lavigne et al. and Clark et al. have compared infection dynamics using analogous PDE and cellular automaton models, and observed qualitatively different infection structures and outcomes between the two approaches [124].
There is a great diversity in the implementation of cellular automata for viral dynamics within the host. Some authors follow a patch–like approach to discretisation of space, such that each spatial region may simultaneously contain many cells (Bernaschi and Castiglione [32], Pandey [157], Rodriguez–Brenes et al. [171], Shapiro et al. [187], for example). While the standard approach under this construction is to keep track of the actual number of cells at lattice sites, in some early approaches, authors have instead considered a binary “high” or “low” concentration of each cell species in each spatial region ( Pandey et al. [157, 156, 158, 142, 135, 136], Kougias and Schulte [120]). As an alternative to patch–like discretisation, spatial regions may represent the cells themselves (Xiao et al. [213], Beauchemin et al. [27], Wodarz et al. [210], Zorzenon dos Santos and Coutinho [224], for example). In these cases, rather than each lattice site tracking a count of cells of different types, each lattice site describes a cell in a certain state which may change depending on whether that cell is susceptible to infection, productively infected, and so on. These cell–regions are often assumed to pack densely and tessellate the entire spatial domain, however some authors have also allowed for empty sites (Wodarz et al. [210, 211], Strain et al. [196], Reis et al. [168]). Bankhead III et al. randomly populated a mostly–empty cell lattice with cell islands whose degree of clustering and confluency was determined from in vitro imaging [22].
There are also highly varied approaches to the representation of viruses in cellular automaton systems. Some authors treat virions, like cells, as discrete agents, and track the movement of individual agents as they spread across the spatial grid (Fachada et al. [67], Shapiro et al. [187, 63], Bernaschi and Castiglione [32], Jafelice et al. [104], for example). Other approaches involve tracking the spread of viral density between lattice sites (Kumberger et al. [123], Immonem et al. [100], Moses et al. [149], for example). Note that we allow such an approach under the cellular automaton classification provided that cells are treated as discrete objects and viral density is described only as a property of the lattice sites. For cellular automata with an explicit viral population, infection is usually implemented (as in the patch framework) via co–location, that is, between virions and susceptible cells which occupy the same lattice site. As an alternative, there is a body of cellular automaton models — particularly in cases where the spatial lattice consists of the cells themselves — which do not consider a free viral population at all (Morselli et al. [147], Jarrah et al. [107], Wodarz et al. [210], Precharattana et al. [165], for example). In such cases, infection is usually implemented not via co–location of infecting and susceptible agents, but by adjacency. For example, this may be encoded as the probability of a cell–region changing state from uninfected to infected, where the probability is dependent on the number of neighbouring cells in an infected state, as in Xiao et al. [213] and Beauchemin et al. [27]. For some, the assumption of neighbour–based infection is considered a proxy for highly diffusion–limited viral spread (Xiao et al. [213], Wodarz et al. [210], for example). Adjacency–based interactions are generally limited to neighbouring cells, although Pandey and collaborators have considered complex mixing operations between lattice sites across the entire grid, in order to represent the migration of cells involved in HIV–host interactions and complex infection spreading dynamics as a result [157, 156, 158, 135, 136, 142].
The discrete nature of cellular automata makes them a useful language for describing stochastic dynamics (although it is worth mentioning that these models are not necessarily stochastic in nature; Zorzenon dos Santos and Coutinho considered a deterministic neighbour–based infection rule [224]). Stochastic dynamics are especially relevant to models at the microscopic scale, such as where the lattice sites are the cells themselves. Wodarz et al. studied infection spreading patterns for oncolytic viruses [210]. Therein, the authors used deterministic approximations of their stochastic cellular automaton model to define regions of parameter space corresponding to different spread patterns, but observed empirically that for parameter values near the boundaries of these regions the cellular automaton the cellular automaton could probabilistically generate different patterns [210]. The authors also found that when the size of the model grid was increased but the initial infection inoculum size is held fixed, there was an increase in the variety of stochastic outcomes of the cellular automaton for the same region of parameter space [210]. Beauchemin et al., by contrast, showed that when the initial infection inoculum is scaled with the increase in size of the spatial grid, it has the effect of reducing stochastic variability of the system [27].
This sensitivity to implementation represents a substantial challenge in applying cellular automata. In addition to the finding by Wodarz and Beauchemin et al. that the dynamics of their models were dependent on the size of the spatial lattice [27, 210], Moses et al. also showed that their SIMCov model generated significantly different dynamics on a two or three dimensional spatial lattice compared to their lung lattice model [149]. The fact of this sensitivity to implementation is not necessarily a limitation of the cellular automaton approach — since it enables the role of structural features in dynamics to be explored — however, this implementation–dependence means that a high degree of care must be taken in approaching the design of cellular automaton models.
4.5 Agent–based modelling: multicellular models
Figure 9: Multicellular models summary. Seminal publications are as follows: Mitchell et al. [145], Yakimovich et al. [215], Eriksen et al. [66], Imle et al. [99], Getz et al. [78], Simmons et al. [190], and Sego et al. [183].
Multicellular models — to follow the nomenclature of Glazier and colleagues [182, 183, 184, 68, 14], for example — are agent–based systems closely related to cellular automata. As with cellular automata, multicellular models involve discrete spatial regions (for example, cells), track discrete numbers of cells of different types, and typically describe meso– or microscopic dynamics. However, the defining feature of the multicellular framework is the presence of multiple spatial scales. These distinct scales typically correspond to the cell scale and a finer–resolution viral scale, for example a discrete cell grid coupled to a continuous virus density surface. There is an important subtlety here: while in practice, such a viral density may be simulated at the cell–scale (that is, in the same manner as a cellular automaton), we consider the model a multicellular system provided that the governing viral dynamics are defined on an alternate spatial scale to the cell lattice (such as by a PDE). This hybrid approach in simultaneously incorporating model components on different length scales has rapidly grown in popularity in recent years among the viral dynamics literature, and is in line with similar approaches in oncology and tissue dynamics, for example [129, 69, 153]. Note that in principle, as we have defined it, this framework could allow for discrete spatial regions to contain groups of cells, however we only found two instances where this was the case. Paiva et al. [155] allowed cancer cells to crowd into a single lattice site; the other model which permitted multiple cells within a single site, by Garrido Zornoza et al. [77], is a phage model defined at ecological scale which is a special case among the models we discuss, as we explore in Section 5.4. All the other models retained in our analysis treated cells as individual spatial regions.
The simplest approach under the multicellular framework is in coupling a cellular automaton to a global virus population. This construction has been used by authors in a variety of applications (Cangelosi et al. [45] and Goyal and Murray [82] in HBV infection, Blahut et al. [37] in HCV, Kreger et al. [121] and Shi et al. [188] in HIV, for example) and generally follows from an assumption that free viruses spread quickly relative to the size of the model tissue [45, 82, 37, 121]. A similar idea was applied by Howat et al., who considered the spread of free virus at the cell scale, but type I interferon (comprised of molecules far smaller than virions which therefore diffuse much more quickly) as a global quantity [97]. Models which consider only infection from a global viral reservoir do not give rise to spatially–structured infections, and as such, the global viral population is incorporated into the model dynamics in different ways. Shi et al. considered neighbour–based infection at a rate related to the global viral load [188]. More typically, however, models with a global virus population generally consider two modes of viral infection: unstructured infection due to the global population of free virions, and structured local infection arising from direct cell–to–cell infection between neighbouring cells (Saeki and Sasaki [173], Kreger et al. [121], Goyal and Murray [82], Blahut et al. [37]). Such models therefore describe two spatial scales of infection, which are generally not well–described by cellular automata. We describe modelling of the cell–to–cell infection route in further detail in Section 6.
A typical approach in multicellular modelling is to embed discrete cells within an underlying continuous spatial domain. Over this domain, authors commonly define the virus population as a density distribution in space governed by PDEs (Mitchell et al. [145], Yakimovich et al. [214], Imle et al. [99], Sego et al. [183], Getz et al. [78], for example). Others include additional small, diffusive species like cytokines or immune components, described in a similar manner (Sego et al. [183], Getz et al. [78], Cai et al. [44], Aristotelous et al. [15] for example). While cellular automata and patch systems cannot easily account for anisotropic spread of virions and other small species, by encoding the dynamics of these components in the language of PDEs, it is natural to account for more complex flows. Yakimovich et al. [214, 215] and Chen et al. [53, 54, 222, 159, 221] added advection to their description of viral density spread and showed the formation of “comet”–shaped plaques (as were observed by, for example, Anekal et al. [13] in their continuum model). Chen et al., in particular, showed that the varied rates of viral advection by the mucociliary escalator in different generations of the lung led to dramatic changes in the spatiotemporal dynamics of infection [53].
This discrete–continuum coupling also enables tracking of the motion of discrete agents. This is of particular relevance in descriptions of the adaptive immune response, which is largely mediated by cytotoxic CD8+ T cells. These cells are relatively few in number and therefore are not well described by a density distribution [126, 111]. As such, Sego et al. [183], Getz et al. [78], Levin et al. [126] and Cai et al. [44] all explicitly modelled discrete CD8+ T cells which migrate towards sites of infection (we discuss immune cell migration in greater depth in Section 6). Some authors have also modelled the individual migration of host cells in the infection of motile cell populations. Imle et al. [99], Bouchnita et al. [40] and Seich Al Basatena [186] all modelled HIV infection among populations of migrating host cells; Jenner et al. [109] also accounted for a spatially–evolving tissue in a model of oncolytic viral infection. In addition to individual cell migration, models which account for the motion of individual virions have also been proposed (Itakura et al. [102], Chen et al. [53, 54, 222, 159, 221]).
One method for implementing this coupled discrete–continuum framework is the Cellular Potts Model [83]. Under this framework, the spatial domain and continuum quantities are discretised over a grid of subcellular voxels. Cells are represented as discrete agents without a fixed shape or lattice: instead, they are characterised by a given target size (in terms of number of voxels), as well as a range of other energy constraints, such as minimisation of boundary length. Over time, the model is simulated by applying stochastic noise at the voxel scale and minimising the energy state of the cell grid. This model therefore accounts for realistic tissue mechanics and packing over time, while providing a natural discretisation for the continuum layer of the model. This framework has been applied to the spatial dynamics of viral infections by Imle et al. [99] and Glazier and colleagues [182, 183, 184, 14, 68]. Getz et al. [78] take a similar approach for viral transmission but use an off lattice point force model for cell mechanics.
The multicellular framework permits the simultaneous representation of multiple biological length scales, which in some ways combines the benefits of continuum and cellular automaton modelling: small particles, especially virions, can be described in terms of density dynamics at the tissue–scale, while the modelling of discrete cells imposes sharp infection boundaries and naturally accounts for cell–scale stochasticity. However, this framework also comes with significant challenges. The substantial leap in model complexity compared to either continuum or lightweight cellular automaton models means that multicellular systems are far more computationally expensive to run. Moreover, this increase in modelling detail is not always accompanied by similarly detailed available biological data. Levin et al., for example, remarked on a lack of data in describing CD8+ T cell motion [126]. This lack of data is especially important since many large multicellular models (Sego et al. [183] and Getz et al. [78], for example) do not (and cannot realistically) calibrate parameter values to biological data using rigorous statistical methods, instead adapting estimated parameter values from elsewhere in the literature. The relevance of such parameter values may vary depending on their context within the model from which they are derived and the biological system to which they are applied.
An additional challenge associated with multicellular systems lies in their implementation. Multicellular models offer a fine resolution of tissues at the cell–scale, however, there is little or no consensus on how such cells should be arranged within the model tissue, for example, or to what extent this influences the dynamics. Some authors use square packing for the cells (Marzban et al. [139, 23] and Chen et al. [54, 53, 15, 222, 159, 221], for example), some use hexagonal packing (Yakimovich et al. [214, 215], Levin et al. [126], Beauchemin et al. [26], Whitman et al. [206], for example), while other consider irregular, time–varying cell packing (Sego et al. [183], Getz et al. [78], Imle et al. [99]). In addition to implementation of the cell sheet, there is also a wide variety in approaches to the representation of continuum quantities relative to the cell grid. We have already discussed the voxel length scale of the Cellular Potts Model; many other authors simply discretise viral and other densities at the cell–scale (Mitchell et al. [145], Beauchemin et al. [26], Holder et al. [95], Howat et al. [97], for example). Others use alternative discretisations independent of the cell grid (Getz et al. [78, 101], Itakura et al. [102], Jenner et al. [109]). This point is important, since, in some models, biological mechanisms occur over small spatial scales, such as cell–to–cell infection between migrating cells (Imle et al. [99], Seich Al Basatena et al. [186]), or viral diffusion in crowded environments (Sego et al. [183], Beauchemin et al. [26], Jenner et al. [110]). Analysis has suggested that in multicellular models of viral dynamics in which virus diffuse sufficiently slowly, the difference between cell–scale and fine–resolution discretisation of the viral density results in qualitative differences in model outcomes [207]. This suggests that the dynamics of such systems are dependent on biological activity at spatial scales far smaller than the cell, which may defy modelling assumptions [207]. Findings such as these point to a need for further discussion and analysis of implementation dependence for multicellular viral dynamics systems.
5 Spatially–structured models by virus species
Much of the specifics of viral infection dynamics can vary substantially between viral species. Different viral pathogens possess unique life cycles, target different cell populations with different biological and structural properties within the host, and have varied interactions with the host immune response. There are therefore a range of modelling particularities which are specific to the viral strain being modelled. This is especially true of spatially–structured models since different host tissues possess specific spatial structures and cells with particular motility properties. In this section, we discuss how spatially–structured models have been applied to the dynamics of specific viral species.
5.1 HIV
Figure 10: Spatially–structured models for HIV. Seminal publications are as follows: Zorzenon dos Santos and Coutinho [224], and Imle et al. [99].
HIV is a chronic infection of lymphoid tissue, particularly CD4+ T cells [160, 151]. Although now well–managed with the daily administration of antiretrovirals, if left untreated HIV infection is characterised by three distinct phases: acute (time scale of days), chronic (time scale of years), and AIDS (time scale of weeks), the last of which is fatal [223, 160, 85]. HIV modelling represents the oldest and most developed branch of the viral dynamics literature in mean–field as well as spatial modelling (reviewed by Perelson [160, 85] and Nowak and May [152], for example), and is the subject of the bulk of early viral dynamics models. In addition to its unique three–phase temporal structure of dynamics, HIV is unusual in terms of the spatial structure of its infection, since its target cells, CD4+ T cells, are motile [85, 99]. Moreover, in addition to releasing virions into the extracellular environment, HIV also spreads largely among host cells through virological synapses formed between host and susceptible cells [118, 39].
The target tissue of HIV infection comprises spatially disparate yet interconnected lymph nodes and lymphatic tissue. Therefore, the definition of model domains and their spatial scale are highly important in modelling HIV. Despite this, many authors define non–specific spatial domains (Zhang et al. [223], Pandey et al. [157, 156, 158, 142, 135, 136], Raza et al. [167], for example). In these cases, the spatial domain might be interpreted as a generalisation of the entire volume of susceptible tissue within the host or alternatively as a representative region of tissue. However, given the structure of the target tissue for HIV, there is a clear difference between these macro– and mesoscopic spatial scales of infection, and such models can therefore only be interpreted as abstractions of the spatial reality. A popular target for many modellers is the within–lymph node dynamics of HIV infection (Bernaschi and Castiglione [33], Zorzenon dos Santos and Coutinho [224], Marinho et al. [137], Strain et al. [196], for example). Though only one section of the overall target tissue is involved in HIV infection, this single–lymph node approach has proved a useful model system for reproducing important global dynamics of HIV such as the three stages of infection (Zorzenon dos Santos and Coutinho [224], Bernaschi and Castiglione [33], Marinho et al. [137]). Jagarapu et al. [105] and Cardozo et al. [48, 47] used lymph node models to describe the dynamics of chronic HIV infection under antiretroviral treatment. They showed that antiretrovirals were heterogeneously distributed in lymph nodes, and in particular, that the drugs were unable to deeply penetrate into lymphoid follicles, facilitating the ongoing replication of virions [105, 48, 47]. We only found one study, by Nakaoka et al., which explicitly modelled the between–lymph node dynamics of HIV infection [151]. In their work, the authors showed that, even for high antiretroviral dosage, there was a limit to the global drug efficacy [151]. This was due to the heterogeneity in drug delivery across the lymphoid network, with some lymphoid compartments not receiving sufficient drug load to inhibit viral replication even when high doses were administered [151]. Other biological domains have also been modelled for HIV. Jafelice et al. [104] and Benyoussef et al. [30] considered HIV dynamics in the blood, while Immonem et al. [100] and Imle et al. [99] modelled HIV dynamics in vitro. These in vitro models serve as useful tools for interpreting biological data, although several groups have noted that it is difficult to build laboratory systems for HIV infection which are faithful to the dynamics within the host, due to the migratory properties of the host CD4+ T cells and the three–dimensional structure of host tissues [99, 103, 118]. Imle et al. applied a detailed interdisciplinary approach to overcome this disconnect by studying the dynamics of a three–dimensional in vitro system which included a complex collagen lattice, designed to simulate the structure of infected tissue within the host [99]. The authors found that these structural considerations hampered the diffusion of free virions compared to other in vitro systems and promoted direct cell–to–cell infection [99].
Authors have approached the motility of the target cells in HIV infection in different ways. There are several cellular automaton approaches to HIV which ignore CD4+ T cell migration altogether and assume static cells and neighbour–based infection (Zorzenon dos Santos and Coutinho [224], Sloot et al. [191], Precharattana et al. [165], for example), the same as has been implemented for viral species which target non–motile cells (Wodarz et al. [210] and Bhatt et al. [36] in oncolytic viruses, Xiao et al. [213] for HBV, for example). Strain et al. applied a similar approach, where instead of being explicitly neighbour–based, infection probability was determined by local virion bursts from static infected cells [196]. While such models have the benefit of simplicity and have been shown to exhibit temporal dynamics which qualitatively agree with clinical and experimental observations of HIV infections, others, including Graw and Perelson, have criticised this approach for its un–biological construction [85]. Despite the high–level agreement with observed dynamics, this disconnect with the biological reality of HIV–infected tissues means that these models should not be interpreted as (spatially) mechanistic. Pandey et al. modelled HIV infection using a static cell grid, but accounted for cell motility implicitly through the rules for infection. The authors used cellular automaton models with a complex (phenomenological) mixing operation between lattice sites to account for experimentally–observed diffuse infection patterns [157, 156, 158, 142, 135, 136]. Other authors have represented the movement of HIV target cells as flux between compartments. In the patch–based approach of Jagarapu et al. [105] and Cardozo et al. [48, 47], CD4+ T cell density had essentially two rates of migration for different regions of the tissue: diffusion–like spread between radial layers of lymph follicles, and instantaneous spread among the (mean–field) free blood/lymph compartment. Bernaschi and Castiglione, by applying the ImmSim model to HIV, tracked the locations of a discrete population of CD4+ T cells as they migrated across a coarse spatial lattice [33, 21, 50]. Seich Al Basatena et al. also tracked the movement of discrete CD4+ T cells across a lattice, but at far higher resolution, such that the cells moved across a lattice of cell–sized, mostly empty sites [186]. More recently, some authors have also modelled the migration of individual T cells in continuous space. Bouchnita et al. [40] and Grebennikov and Bocharov [89] considered spherical cells whose displacement was governed by Newtonian motion. At an even finer level of spatial detail, Imle et al. [99] also used physics–based rules for CD4+ T cell migration, but also allowed for cellular deformation (through modelling in the Cellular Potts Model framework, discussed in Section 4.5).
5.2 Oncolytic viruses
Figure 11: Spatially–structured models for oncolytic viruses. Seminal publications are as follows: Wu et al. [212], Wein et al. [205], Friedman et al. [73], Mok et al. [146], Wodarz et al. [210], Malinzi et al. [132], Friedman and Lai [72], and Alzahrani et al. [9].
Oncolytic viruses are a broad group of genetically–engineered viruses that preferentially infect, replicate within, and kill tumour cells, while leaving healthy cells untouched. This, in principle, makes the prospect of oncolytic virotherapy an attractive treatment option for cancers, although experimental and clinical results have so far been mixed, and development is ongoing [125, 108]. The complex, multiscale dynamics of oncolytic viral infection, as well as the marked spatial constraints within tumours, makes oncolytic viruses well suited to spatially–explicit mathematical modelling, although such viruses are a special case among the viral dynamics literature. For one, since oncolytic viral infections are deliberately introduced into host tissue, the placement and dosage of the viral inoculum are especially important [205, 212, 20]. Moreover, since infection takes place in pathologically crowded tissue under high pressure, host tissue is subject to constant movement and rearrangement, and free viral diffusion is highly constrained [109, 147].
Spatially–structured models for oncolytic viruses are overwhelmingly implemented as PDE systems. There are several reasons for this. For one, it is clearly of interest to model virus–host interaction at the whole–tumour scale, which is generally well approximated by a density description for the model components. PDE systems also allow description of the mechanics of the intra–tumour density (Morselli et al. [147], Friedman and Lai [72], Friedman et al. [73]) and the tumour–healthy tissue interface (Alzahrani et al. [9, 10], Alsisi et al. [6, 7, 5]). In simpler models, the growth of the tumour is approximated as radially symmetric, reflecting the approximately spherical structure of solid tumours in the early stages of growth (Malinzi et al. [132], Friedman et al. [73], Wu et al. [212], Wein et al. [205], Mok et al. [146]). This assumption reduces the model to one–dimensional in space, which simplifies mathematical analysis. The aims of analysis of models of oncolytic viral dynamics are often different to those in analysis of the dynamics of other viruses, and more often focus on the tumour, rather than the infection. While for, say, SARS–CoV–2, authors have derived approximations of the speed of the infection front (Tokarev et al. [201], Ait Mahiout et al. [2, 1]), in models of oncolytic viral infections, estimates have instead been made for the speed of the tumour front (Malinzi et al. [132], de Rioja et al. [60], Wein et al. [205]). Wu et al. [212] and Wein et al. [205] also used analytical techniques to derive expressions for tumour extinction thresholds as functions of the model parameters.
A popular application of models for oncolytic viral infections is in testing the impact of different approaches to viral administration to the tumour. Wu et al. [212] and Wein et al. [205] both studied models for radially symmetric tumours with virus initially delivered either to the rim or to the core of the tumour. Wu et al. suggested that, taking into account the necrosis of the tumour core due to nutrient diffusion from the rim, core injection was likely to be somewhat redundant, and that rim injection would be more effective in spreading through the tumour [212]. Wein et al. found that for oncolytic virotherapy to be effective, the virus must be highly aggressive — that is, extremely infectious and fast–spreading — in order to outpace the growth of the tumour [205]; this result was recapitulated by Al-Johani et al. [4] and Simbawa et al. [189]. Pooladvand et al., however, has warned that high infectivity is not necessarily sufficient for complete eradication of the tumour [164]. Other authors have also found that the success of oncolytic virotherapy depends on viral infection reaching widespread regions of the tumour. Storey and Jackson compared a viral dosage strategy which repeatedly targeted the same location in the tumour to an adaptive strategy which instead delivered oncolytic viral load to the location of highest tumour cell density at that point in time, and found that the adaptive strategy dramatically reduced the final tumour mass after the same length of treatment time [195]. Bailey et al. proposed a model with randomly seeded foci arising from viral extravasation but with a hard limit on the radius of infection foci, following experimental observations of limited focal spread [20]. This approach suggested that, in order for virotherapy to be effective, the number of viral foci would have to be extremely high, such that the overall infected volume would cover a sufficiently wide region [20].
This spatially constrained spread of infection foci within tumours is a recurring theme in explaining the limited in vivo efficacy of oncolytic viruses. One reason for this limitation in the spatial spread of infection is due to the fact that the extremely dense tumour environment severely inhibits the diffusion of free virions [110, 147]. As a result, many authors have argued that infection is likely only possible directly between neighbouring cells (Morselli et al. [147], Bhatt et al. [36], Reis et al. [168]). Wodarz et al. showed that experimentally–observed infection patterns for oncolytic viruses could be generated from a cellular automaton with adjacency–based infection only [210]. As a result, it has been widely argued that increased diffusion of free virions would dramatically improve oncolytic virotherapy (Paiva et al. [155], Jenner et al. [110, 109], Mok et al. [146], for example), however, there is disagreement in the modelling community about the conditions under which this might come about. Mok et al. [146] and Paiva et al. [155] have argued that a reduction in density of the tumour extracellular matrix would increase the rate of viral diffusion, however, Alsisi et al. [6, 7, 5] have suggested that virions preferentially migrate towards regions of higher matrix density via haptotaxis. A related problem to the rate of viral diffusion is the relative density of the tumour cells and normal stromal cells. Since the latter cannot be infected by the oncolytic virus, Wein et al. [205] and Jenner et al. [110] have shown that oncolytic viral infections are far more effective in tumours where cancer cells are densely packed, with few interspersed stromal cells. However, in contrast to these findings, Berg et al. showed that oncolytic infections spread more slowly in a three–dimensional tumour model compared to an equivalent model in two dimensions, despite cells in the former having a higher number of adjacent tumour cells [31].
Since the spread of viral infections in tumours is mainly between neighbouring cells, the arrangement of cells in space is highly important in agent–based modelling of oncolytic viruses. Simpler approaches to the arrangement of tumour cells in space consider a regular lattice of cell–sites which remain fixed over time (Wodarz et al. [210], Reis et al. [168], Paiva et al. [155] for example). Others have constructed spatially–static networks of cells based on Voronoi tessellations derived from experimental imaging of tumours, which provides a realistic distribution of the number of neighbours between cells across the tumour (Berg et al. [31] and Bhatt et al. [36]). Berg et al. compared oncolytic viral dynamics on a regular and Voronoi cell lattice in three dimensions and found that infections on the Voronoi lattice were less likely to result in eradication of the tumour [31]. Temporal evolution of the cell grid has also been implemented. Jenner et al. considered a time–varying Voronoi lattice model for tumour cells, which also enabled the explicit representation of tumour cell proliferation and remodelling, and moreover facilitated the simulation of time–varying adjacency between cells [109].
5.3 Respiratory viruses
Figure 12: Spatially–structured models for influenza and SARS–CoV–2. Seminal publications are as follows: Beauchemin et al. [27], Mitchell et al. [145], Getz et al. [78], Quirouette et al. [166], and Sego et al. [183].
Viral infections that target the respiratory tract — including influenza and SARS–CoV–2 — have had a spike in popularity among the modelling literature following the outbreak of the COVID–19 pandemic. Respiratory viruses, unlike HIV or oncolytic viruses, broadly follow the dynamics of the “general” viral infection as modelled in the literature: that is, respiratory viral infections spread among non–motile cells within a single, contiguous body of tissue, across which cell–free virions can diffuse relatively freely [76, 26, 95]. Nonetheless, there are still elements of spatial infection dynamics which are specific to infections of the respiratory tract. We note that, to the level of detail that mathematical models have described them, there is little difference between the dynamics of different respiratory viral species; as such in the following we discuss models for respiratory viral infections in general.
Viruses which infect the respiratory tract specifically target the monolayer of epithelial cells which line the airways [53, 149, 76]. Thus, to an extent, the natural geometry of respiratory infection is two–dimensional, especially at the microscopic scale, and as a consequence virtually all of the respiratory infection models retained for analysis from our search also consider two spatial dimensions. However, at a larger length scale, there is substantial heterogeneity in the structure of the lung along the length of the respiratory tract. The respiratory tract has a complex, tree–like structure, comprising 24 generations of branching airways, such that upper passages like the trachea are relatively wide (on the order of centimetres), while deeper airways are very narrow (on the order of fractions of millimetres) [130, 53]. Ayadi et al. [16] and Moses et al. [149] have explicitly considered the dynamics of viral infection across the entire, heterogeneous lung. In the latter case, Moses et al. constructed a full–size, three–dimensional digital human lung, complete with a detailed branching structure for the cells [149]. The authors showed that viral dynamics on this highly complex structure were significantly different to the dynamics of the same infection model on a purely two–dimensional cell grid [149]. While the use of this hyperrealistic geometry demonstrates the notable differences in the dynamics of infections on flat sheets compared to the lung, there is a substantial conceptual leap between the two geometries. Other authors have instead employed an intermediate approach, studying the spread of viral infections on sections of the lung in order to identify specific ways in which the structure and geometry of lung tissue influence the spread of infection. Chen et al. [53] and Williams et al. [208] studied the spread of viral infections on individual lung branches, or on small sections of branching tissue. The latter group showed that tubular and branching tissue geometry imposes rich, spatially–directional infection dynamics [208]. These dynamics moreover exhibited an increased sensitivity to the rate of viral diffusion compared to the dynamics of infection on a flat cell sheet [208].
In addition to structural heterogeneity, there is also variability in the cellular composition of upper and lower regions of the respiratory tract. As a consequence, the rate of spread for different viral strains varies for different regions of the lung depending on their cellular tropism [74, 76]. Tokarev et al. [201] and Ait Mahiout et al. [2] calibrated PDE models to experimental data and used travelling wave methods to compare the dynamics of the omicron and delta variants of SARS–CoV–2. The authors found that the omicron variant had a more rapid invasion speed and therefore a fitness advantage in the early stages of infection [201] and in the upper regions of the respiratory tract [2]. However, Tokarev et al. also found that in the delta variant spread faster in the lower respiratory tract [201]. These findings are consistent with the usual experimental and clinical observation that viruses which spread preferentially in the lower respiratory tract (like avian–adapted influenza strains or the delta variant of SARS–CoV–2) are typically associated with high pathogenicity, while those that spread preferentially in the upper respiratory tract (like seasonal influenza strains or the omicron variant of SARS–CoV–2) are typically more infectious but result in milder disease [192, 76, 112, 74].
Another characteristic feature of the lung is the mucociliary escalator, the directed flow of the mucus lining the respiratory epithelium from lower to upper regions of the respiratory tract. Quirouette et al. [166] included the mucociliary escalator in a model of influenza infection, and suggested that infection can effectively only spread up the respiratory tract from the initial point of virion deposition [166]. The rate of mucociliary flow is variable at different depths of the respiratory tract, with slower flow in deeper regions of the lung [130, 53]. Chen et al. compared the dynamics of SARS–CoV–2 infection in airway tissue from different generations of the respiratory tract and suggested that infection spread was severely limited in the lower airways [53]. The authors postulated that infection propagation in the deep lung must therefore be driven by a large quantity of viral depositions from upper regions of the lung, such as virus–laden droplets or aerosols [53]. Chakravarty et al. — who extended the Quirouette model — reached a similar conclusion, arguing that aerosolisation of infected nasopharyngeal mucosa (ANM) was necessary for infection to reach the deep lung [52]. Vimalajeewa et al., by contrast, modelled advection down the respiratory tract in the course of SARS–CoV–2 infection due to air flow in the respiratory lumen (the authors accounted for the mucociliary escalator as a drag force and slip correction factor in the flow model) [203]. In addition to its role in disease progression, Zhang et al. suggested that the rate of mucus advection also influences the nasal viral load [221]. The authors argue that variation in this rate between individuals could offer one potential explanation for the extreme variation observed in clinical nasal swab titres among COVID–19 patients [221].
5.4 Other viruses
Figure 13: Spatially–structured models for other specific viruses. Seminal publications are as follows: Xiao et al. [213], Heilmann et al. (2010) [93], Heilmann et al. (2012) [94], Schiffer et al. [177], Eriksen et al. [66], and Simmons et al. [190].
In addition to the viruses which we have discussed above, several other species have been considered. Among these, a substantive body of work has been developed for the modelling of bacteriophage (or simply “phage”) dynamics. Phage and the bacteria which they infect are both some of the largest biological populations in existence, and their interaction is one of the oldest predator–prey relationships on Earth [190, 144]. Bacteria and phage inhabit and interact within a broad array of environments, including in soil, water, and within larger organisms [93, 77]. We note that there is a major difference between phage dynamics and the dynamics of the other virus species which we have considered in this review, in that phage hosts are single–cell organisms. As such, while models for phage–bacteria interactions describe the spread of infection among cells — as with the other models we have discussed — such systems are, strictly, between–host models. We have nonetheless included such studies in our analysis.
Due to the unique properties of phage–bacteria dynamics, as compared to the dynamics of other viral species we have discussed, modellers have considered the spatially–structured interactions between phage and microbe communities on far larger spatiotemporal scales. Garrido Zornoza et al. modelled the spread of phage among a microbial community at an ecological spatial scale (the only such model we found in our search), which described spatially explicit microbial habitats, from which both viruses and host microbes were subject to stochastic aerosolisation and advection with wind [77]. The authors showed that this influx–outflux process among habitats lead to the spatial spread of microbial communities but also resulted in destabilisation and occasional local extinctions in individual habitats [77]. While phage–bacteria systems exhibit spatially–structured dynamics on this ecological spatial scale, most authors instead model the dynamics of phage–bacteria interactions on far smaller length scales. In particular, many authors have modelled the spread of phage infection among dense, surface–bound microbial communities called biofilms. The crowded nature of these environments requires particular consideration by modellers. Martínez-Calvo et al., for instance, considered density-dependent motility of bacteria and showed that the spread of phage infection among the colony can disrupt bacterial aggregation [138]. Moreover, as with oncolytic viral infections in tumours, the densely–packed environment of biofilms almost entirely suppress viral diffusion, resulting in highly localised viral spread. Taylor et al., for example, showed that the high cell density found in biofilms promotes the formation of spatial clustering of infection [198]. Modellers have accounted for the crowded nature of biofilms in various ways. While some, including Heilmann et al. (2012) [94] and Mitarai et al. [144], have implemented linear diffusion to describe the dispersal of cell–free phage with a very small diffusion coefficient compared to say, the Stokes–Einstein estimate (Akpinar et al. [3], Beauchemin et al. [26], Sego et al. [183], for example, took similar approaches for other viral species). Other authors have instead opted for non–linear forms of the diffusion term: Hunter et al., for example, considered density–dependent diffusion [98]; Simmons et al. explicitly considered physical blocking of phage motion by bacteria [190]. Valdez et al. showed that, as a consequence of this highly localised spread, differently–shaped initial distributions of phage among bacterial communities gave rise to spatially distinct colony structures [202].
One of the remarkable features of phage–bacteria interactions is the coexistence of both populations over vast time scales, even within a single biofilm. This relationship is a property of the rich and finely–calibrated predator–prey dynamic between phage and host microbes. Several authors have shown that spatial structure in phage distribution within microbial communities is a key contributor to long–term stability of both populations. Heilmann et al. (2010) compared coexistence of phage and bacterial populations in a cellular automaton model with either spatially–localised or well–mixed conditions, and found that the former permitted coexistence over a far larger region of parameter space [93]. Authors have proposed several mechanisms by which spatial structure permits coexistence. In their study, Heilmann et al. (2010) allowed infection of dead or already–infected bacteria by phage, and demonstrated a “shielding” effect, whereby infected or dead regions of the biofilm would act as phage sinks, protecting regions of susceptible bacteria [93]. In a different study, the same group (Heilmann et al. (2012)) [94], demonstrated the spontaneous formation of these spatial bacterial “refuges” in their models, a phenomenon also shown by Bull et al. [42]. Others have shown that the spatial structure of infection influences the complex co-evolutionary dynamics between phage and bacteria. Haerter and Sneppen modelled the evolution of phage resistance mediated by the Clustered Regularly–Interspaced Short Palindromic Repeats (CRISPR) mechanism and showed that well-mixed conditions resulted in inefficient and Darwinian evolution of resistance, while spatially–structured conditions resulted in Lamarckian evolution and provided better agreement with experimental evidence [91]. Others have sought to explain the counter-intuitive experimental observation that phage appear to be able to develop virulence faster than bacteria can develop defences, yet do not drive host populations to extinction [93, 94]. Coberly et al. suggested that in competition between moderate and highly–virulent phages, the high selection pressure for resistance to the virulent strain would favour the persistence of the weaker virus, despite the moderate virus being otherwise outcompeted [55]. Notably, the authors suggested this was a direct consequence of the spatial structure of infection, since they observed “percolating clusters” of resistant bacteria which provided spatial refuges for the moderate phage [55]. In another study, Eriksen et al. suggested that the survival of phage populations among spatially–isolated bacterial microcolonies relies on the phage sustainably “farming” colonies instead of destroying them [66].
Hepatitis B and C viruses, which cause chronic infections of the liver, have also been modelled by several authors. Despite the recent development of highly effective direct–acting antivirals for the latter, hepatitis viruses are still globally widespread with relatively high rates of mortality [90, 58]. Xiao et al. [213] and Cangelosi et al. [45] have argued that spatial heterogeneity of infection may play a role in immune evasion, supporting viral persistence in the liver. Goyal and Murray [82] and Blahut et al. [37] have also argued that direct cell–to–cell transmission of hepatitis viruses also contributes to the chronicity of infection (we discuss modelling for cell–to–cell infection in greater depth in Section 6). Moreover, within–cell dynamics appears to be key in hepatitis B dynamics. Cangelosi et al. showed that viral cccDNA represents one of the most robust obstacles to hepatitis B treatment, and suggested that even a single cccDNA molecule would be sufficient to reinfect the entire liver [45].
Some authors have also modelled the spatial dynamics of herpes simplex virus (HSV–2). HSV–2 infections target the genital tract and are also chronic [181]. Schiffer et al. showed that, during viral shedding episodes, HSV–2 replicates extraordinarily quickly, and moreover rapidly spreads in space, effectively “outrunning” the initial immune response and prolonging the episode [177]. Schiffer [176] and Dhankani et al. [61] have also highlighted the highly complex and spatially–structured interaction between HSV–2 infections and host immune reponses. Dhankani et al. suggested that while the dynamics of HSV–2 episodes could be partially predicted in the short term based on the spatial distribution of immune cell density, on the time scale of weeks or months the dynamics of the episode became impossible to predict [61]. HSV–2 infections have also been modelled in co–infection with HIV. Modelling has shown a synergistic relationship between the two species: Byrne et al. suggested that HSV–2–infected women were up to 8.6 times more likely to acquire HIV infection than healthy hosts [43]. On the other hand, Schiffer et al. showed that, as a result of CD4+ T cell depletion, HIV–infected individuals were far more susceptible to HSV–2 infection [180].
We also found a single publication which modelled plant viruses (Rodrigo et al. [170]), along with a handful of models for Epstein–Barr virus (EBV). EBV is an extremely common viral infection in humans which causes generally mild illness but remains dormant in hosts and may be reactivated, occasionally causing long term complications [187, 63]; it has also been associated with the development of cancers [49]. EBV targets the tonsils of Waldeyer’s ring: Shapiro et al. used a cellular automaton model which explicitly accounted for the entire anatomy of Waldeyer’s ring and its peripheral circulation [187, 63]. In a separate model, Castiglione et al. showed that the persistence of EBV in the host was due to latently–infected B cells spreading to this peripheral pool, where they were not subject to immune surveillance [49].
6 Spatially–structured dynamics
Mean–field models, especially ODE models, have a long history of application to the dynamics of viral infections and have offered substantial insight (reviewed in [160, 193, 152, 192], for example). However, there are a range of aspects of viral infection dynamics which inherently involve a spatial mode of action, such as the action of the immune system, or the spread of infection in crowded tissue. The mismatch between the spatially–structured biological reality and these non–spatial models means that there is a limit to the precision possible with such models in representing accurate infection dynamics, or in accurately determining parameter values [24, 76]. In this section, we review how spatially–structured models have accounted for spatially–acting aspects of viral infection, and discuss how these dynamics differ from descriptions by non–spatial models.
6.1 Spatial heterogeneity and parameter estimation
One of the simplest improvements offered by spatially–structured modelling is in allowing for spatially–heterogeneous distribution of infection. We have already discussed several examples of heterogeneous tissues, including heterogeneity in pressure and density within tumours for oncolytic viruses (Morselli et al. [147], Alsisi et al. [6, 7, 5]), variable viral advection rate at different depths in the lung due to the mucociliary escalator (et al. [53, 222, 54, 159], Vimalajeewa et al. [203]), and structural heterogeneity in the arrangement of cells (Bhatt et al. [36], Jenner et al. [109], Berg et al. [31]). Heterogeneity of tissue has been shown to qualitatively impact the outcome of models compared to homogeneous systems. Chen et al. showed that the probability of the same viral inoculum establishing infection varied dramatically at different depths of the respiratory tract, based solely on variations in the rate of mucociliary flow [53]. In a patch model for viral dynamics, Funk et al. implemented random heterogeneity of model parameters between patches and found that this heterogeneity impacted the dynamics of the equilibrium state of infection [75].
Beyond these qualitative differences in infection patterns, spatially–varying infections have been shown to have a profound quantitative impact on parameter estimation. Funk et al. also showed that in one region of parameter space, the patch model predicted viral persistence while an analogous mean–field model predicted extinction [75]. More explicitly, Strain et al. compared simulations of HIV dynamics using both a cellular automaton and an ODE system, and found that the mean–field model overestimated viral infectiousness more than 10–fold [196]. Mitchell et al. obtained a similar result in influenza [145]. In their study, the authors calibrated both a cellular automaton and an ODE system to experimental data, and found that the predicted basic reproduction number, $\mathcal{R}_{0}$, for pandemic–strain influenza was estimated about 5–10 times higher using the ODE model [145]. Clearly, this has dramatic consequences for the characterisation of the virus.
6.2 Cell–to–cell viral transmission
ODE models for viral dynamics rely on an assumption that cells and virions are well–mixed. Under this construction, virions have equal access to all cells, and all uninfected cells are equally likely to become infected, regardless of their position in space. Under the standard assumption that infection is propagated through the dispersal of cell–free virions, this corresponds to an assumption that virions spread quickly across the model tissue (Holder et al. [95], Beauchemin [25]); we have already discussed some instances where this assumption might be appropriate (Cangelosi et al. [45], Blahut et al. [37], Kreger et al. [121], for example). However, there has been growing recognition, both in the biological and modelling literature, of the importance of the cell–to–cell route of infection [106, 122, 119, 200, 76]. This mode of transmission has been observed in infections by virtually all medically–important viruses, including retroviruses, herpesviruses, coronaviruses and influenza viruses [200, 175]. Since cell–to–cell transmission relies on adjacency between cells, spatial structure is a central aspect of its mode of action. Crucially, although this phenomenon has been described using ODE methods, this approach has had limited success. Kumberger et al. generated synthetic time series data of viral infection using a cellular automaton which accounted for both modes of infection, then attempted to refit the data using ODEs [123]. The authors found that ODE models could not accurately describe the dynamics unless additional terms were included to artificially account for the spatial structure [123]. A similar conclusion was reached by Graw and colleagues in an ODE modelling study (not included in our search since this was a non–spatial model), who attempted to fit experimental data on cell–to–cell infection and also needed to artificially account for the spatial structure of the infection to obtain a reasonable fit [84]. Williams et al. conducted simulation–estimation studies on both ODE and multicellular models for viral dynamics with both modes of infection and found that the balance of the two modes of spread in the ODE model was not practically identifiable in the presence of observational noise [209]. This finding suggests that the two mechanisms are relatively interchangeable in these non–spatial models [209].
Spatially–structured modelling has provided insight in characterising the role of the two modes of infection in overall viral infection dynamics. Blahut et al. fit a multicellular model to experimental data for hepatitis C and concluded that cell–to–cell infection could account for as much as $99$% of infection events, however cell–free infection was crucial to the overall progression of the virus by providing access to uninfected regions of the tissue away from the infection focus [37]. Goyal and Murray, by contrast, argued that cell–to–cell infection was not especially important to the overall dynamics of hepatitis B in the early stages of infection, but was crucial in the chronic stage of infection through T cell exhaustion [82]. On this longer time scale, Saeki and Sasaki [173] and Kreger et al. [121] have also implicated cell–to–cell infection in the evolutionary dynamics of viral infections (we discuss evolutionary dynamics in greater depth in Section 6.3).
However, while modelling has offered qualitative insights into the role of cell–to–cell transmission, there remains substantial quantitative uncertainty as to the actual amount of cell–to–cell transmission (relative to cell–free infection) in the spread of different viral infections. This represents a substantial modelling challenge for several reasons: for one, this quantity is likely to be substantially different for different viral species [220, 37, 119, 99], and moreover, the balance of the two roles of transmission is expected to vary between in vitro and in vivo contexts [62, 65]. Blahut et al. [37] and Durso–Cain et al. [65] estimated the relative contribution of the two modes of viral spread by comparing experimental data with and without the presence of extracellular virus–neutralising antibody. This approach, however, is limited to in vitro settings, and incurs additional complexity by introducing the possibility of inhibition or enhancement of the cell–to–cell route in the presence of cell–free infection blockage. Kumberger et al. [123] and Williams et al. [209], by contrast, each attempted to estimate the relative proportion of the two modes of infection using only a single simulated dataset, describing uninhibited viral dynamics. Both groups determined that global infection metrics alone were insufficient to practically identify the balance of the two modes of infection, however, Williams et al. [209] showed that, in principle, data describing the spatial structure of the infection would be sufficient to enable this quantity to be inferred.
Due to the migration of target cells in space, the role of cell–to–cell transmission in HIV is somewhat unique. A detailed discussion was provided by Imle et al. [99]. The authors simulated HIV dynamics using a Cellular Potts Model calibrated to detailed imaging data to track the movement and contact between CD4+ T cells. Upon comparison with experimental time series data, the authors concluded that cell–to–cell transmission of HIV requires at least $25$ minutes of contact between donor and recipient cells [99]. The authors also modelled in vitro data collected from cells either in suspension or in a 3D collagen matrix — which was argued to be representative of conditions in vivo — and found that cell–to–cell infection was significantly promoted in the collagen matrix, compared to the suspension [99]. However, this calculation was derived using fits to non–spatial ODE models, and, perhaps as a consequence, estimates for the prevalence of the cell–to–cell mechanism were subject to extremely wide uncertainty [99].
6.3 Spatial immune activity
In addition to the spatial spread of infection, there are also aspects of the immune response which are inherently spatially–structured. Modelling has shown that the presence of spatial structure significantly increases the potency of the interferon–mediated innate immune response (Lavigne et al. [124], Segredo–Otero and Sanjuán [185]). Lavigne et al. modelled innate immune response to viral infection using ODE, PDE and cellular automaton techniques [124]. The authors demonstrated the formation of buffer regions comprised of interferon–mediated antiviral cells surrounding the infection focus, which they likened to a “ring vaccination” strategy as in epidemiology [124]. The cellular automaton model, owing to its stochasticity, also facilitated the possibility of temporary viral resurgence, due to rare long–range infection events beyond the protective region [124].
In the adaptive immune response, T cells are recruited from the lymph nodes, then traverse infected tissue along chemotactic gradients to the site of infected cells. This search process has been described by a number of authors (Beauchemin et al. [27], Sego et al. [183], Getz et al. [78], for example). Levin and colleagues used a multicellular model to characterise the dynamics of the T cell search in the adaptive immune response to influenza A infection, and showed that the window of time which T cells have to locate and clear infection foci is dependent on the viral strain [126]. A similar conclusion was reached by Morselli et al. in the context of oncolytic viral dynamics, who suggested that if the adaptive immune response is activated too quickly following the initiation of virotherapy, the infection has limited effect on the growth of the tumour [148]. Levin et al. moreover suggested that the ability of the immune response to effectively clear infection relies on viral diffusion being sufficiently slow: beyond a certain threshold, the infection spreads too quickly for the T cell response to keep up [126]. This serves to highlight the importance of space in the action of the immune response to viral infections. Kadolsky and Yates, too, showed that under the assumption of chemotactic T cell migration, the infection structure in space is crucial in determining the outcome of the immune response, with highly clustered infections significantly more likely to be cleared [111]. As a consequence, since infection typically spreads patchily within the host, the authors predicted that mean–field models would overestimate the number of T cells needed to clear infection [111].
Just as spatial structure plays a key role in immune action, it is also important in the evolution of immune–resistant traits among the viral population. Capitán et al. [46] and Lavigne et al. [124] have demonstrated that spatial confinement of infection dramatically increases evolutionary pressure on viruses, and Phan and Wodarz [161] showed that spatial constraints promoted competitive exclusion between two competing viral strains and increased the likelihood of one strain becoming extinct. Yin [216] showed that spreading of viral plaques in space promotes the development and spatial separation of viral mutants. Segredo–Otero and Sanjuán discussed the role of spatial structure in the evolution of evasion of the innate immune response specifically [185]. The authors showed that in infection with a mix of interferon–triggering and interferon–evading viral variants, the immune–evading population are selected for and the immune–triggering variants are eliminated only when spatial structure is present [185]. Sardanyés et al. also suggested that spatial structure selects for viral genotypes which are robust against deleterious mutation (so–called “survival of the flattest”) [174]. Spatially–structured immune dynamics have also been used to explain evolutionary strategies of both viruses and host immune responses. Komarova used a model of viral–antibody interaction to show an evolutionary advantage of lytic replication of viruses, that is, virions which replicate inside host cells but are only released by killing (“bursting”) the cell [117]. The model demonstrated a “flooding” effect, where the bulk release of virus in bursting events allowed virions to spread far enough, without being neutralised by antibodies, to infect new cells and continue the infection [117]. By contrast, the model suggested that a steady, budding release of virions would be more effectively controlled by the antibody response [117]. Cuesta et al. probed this phenomenon further in the context of evolutionary competition between virus and host, for instance in phage–bacteria dynamics [57]. The authors determined that a large burst size is advantageous for the virus only when host cells are readily available: in spatially–restricted settings this strategy offers no such advantage [57]. Seich Al Basatena et al. studied the role of lytic and non–lytic CD8+ T cells in viral escape during HIV infection, and found that while a lytic CD8+ T cell response rapidly selected for an escape variant virus, even a non–lytic response (in the sense of secreting antiviral chemokines that on contact with susceptible cells block any infection) was found to promote viral escape [186]. This effect was due to the spatial structure of infection, since both wild–type and variant viruses were observed to form infections in clusters, only the former of which could be recognised by the CD8+ T cells [186]. As a consequence, the resulting distribution of antiviral chemokines were observed to spatially coincide with wild–type infected regions, while leaving variant–infected regions less likely to be affected [186].
Saeki and Sasaki [173] and Kreger et al. [121] have also discussed the role of the cell–to–cell and cell–free mechanisms of infection in evolutionary dynamics. Saeki and Sasaki used Monte Carlo simulation to predict the evolution of a preference for either mechanism where infection took place on a cell grid, a given proportion of which were selected to be susceptible [173]. The authors showed that cell–to–cell spread tended to be selected for the more that cells were clustered together [173]. Conversely, Kreger et al. considered the evolution of HIV under varying proportions of infections arising from the cell–to–cell route [121]. The authors found that a cell–to–cell spread was conducive to the prevalence of disadvantageous mutants, which may facilitate the pre–existence of drug–resistant variants within the host [121].
7 Future development for modelling of spatial dynamics of viral infections
Spatial structure has come to be recognised as an important factor in the dynamics of viral infections. Here, we have conducted a scoping review to show the huge diversity in approaches to modelling spatial aspects of viral infections. The wide range of methods and approaches used mirrors the significant diversity in modelling requirements between applications, including particularities of viral species or model tissues, for example. We have shown that spatial models offer a description of a range of biological phenomena with an intrinsically spatial mode of action, such as cell–to–cell transmission, where an accurate description is not possible using traditional ODE modelling. Spatial models have moreover been shown to demonstrate substantial differences in parameter estimates compared to non–spatial approaches, suggesting that this approach to modelling may be necessary to improve quantitative biological insights as well. However, there remain important challenges to be addressed if spatially–structured models are to become more widely used in biological applications.
7.1 Robust implementation
One of the key observations from our scoping review is the presence of an incredibly wide array of approaches to representing spatial structure in viral infections. While it is natural — indeed, essential — that there is debate among the modelling community over which components and interactions should be captured in a model, this lack of consensus is particularly prevalent among spatial models and at a much more fundamental level. While traditional, non–spatial models for viral dynamics more or less share the common language of ODEs, there are fundamental differences between, say, PDE models of viral dynamics and cellular automata. Our analysis has shown that this diversity in approaches is not a shortcoming of the field, but a necessity in order to describe the highly varied spatial dynamics of different viruses at different scales. However, future models should be developed with a clearer sense of the scope and limitation of their approach. PDE models, for instance, provide a good macroscopic description of tissue, yet due to a lack of discreteness or stochasticity, should not be relied upon to account for rare events such as extinction or long–range viral dispersal. On the other hand, cellular automata (perhaps with the exception of the whole–lung model by Moses et al. [149]) should be considered microscopic models only of the small patch of cells which they comprise and — depending on the structure and arrangement of the overall tissue — not necessarily a description of dynamics on the macroscopic scale.
We also showed that a substantial challenge in using agent–based models in particular is a dependence of the dynamics on the exact implementation of the model. As we discussed in Section 4, Wodarz et al. [210] and Beauchemin et al. [27] have both shown that stochastic properties of cellular automaton models are influenced by the size of the grid. Berg et al. [31] and Moses et al. [149] have also demonstrated that the dynamics of their models varied significantly depending on whether two or three spatial dimensions were considered. This finding is not necessarily a limitation of the agent–based approach: the dependence of model dynamics on dimensionality, for example, is an important biological finding in its own right. Nonetheless, this implementation–dependence clearly highlights the importance of biological fidelity in model construction, and moreover points to a need for careful study of various elements of the construction of agent–based models and how they might influence the dynamics. Key questions, such as the spatial arrangement of cells, the size and dimensionality of the spatial domain, or the discretisation of computational multicellular systems, have so far been largely unexplored in viral dynamical models, yet are crucial for ensuring robust implementation of agent–based systems.
7.2 Application to experimental data
Perhaps the greatest obstacle to the wider usage of spatially–structured models of viral dynamics is the difficulty in calibrating such models to experimental data. Spatial models are inherently more complex and computationally expensive than their non–spatial counterparts and therefore both obtaining sufficient spatial data and carrying out rigorous parameter estimation presents something of a challenge [85, 86, 24, 76]. While some authors have conducted such a fitting process using spatial models (Holder et al. [96], Haseltine et al. [92]), the vast majority instead rely on drawing at least some parameter estimates from elsewhere in the literature. However, the exact meaning of a particular parameter is likely to vary somewhat between models depending on what components and interactions are included. For example, parameter estimations between a model with an explicit eclipse phase and a model without one predict the same time interval between infection generations, but will predict different life spans for infected cells due to the latently infected period [19]. This parameter unreliability is especially notable when parameter estimates for spatial models are obtained using ODE models. We have already discussed examples of spatial and non–spatial models generating wildly different parameter values for the same overall dynamics (Strain et al. [196], Mitchell et al. [145]). As such, parameter estimates drawn from external sources cannot necessarily be treated as reliable for a given alternative application.
One parameter that is almost always cited from other literature is the diffusion coefficient of extracellular virus. In many cases, this coefficient describes the rate of the only spatially–explicit phenomenon of the model (Mitchell et al. [145], Beauchemin et al. [26], Marzban et al. [139], for example), and therefore is by far the most difficult parameter to estimate. The viral diffusion coefficient is usually estimated using the Stokes–Einstein relation (Strain et al. [196], Beauchemin et al. [28], Sego et al. [183], for example). For influenza viruses in plasma at body temperature, the diffusion coefficient has been calculated at approximately $3.18\times 10^{-12}\text{m}^{2}\text{s}^{-1}$ [26]. However, critically, this value is usually not the one actually used in simulations. Instead, authors generally use a value some orders of magnitude smaller. There are varying justifications for this. Akpinar et al. [3] used a smaller diffusion coefficient to obtain qualitative agreement with experimentally–observed spatial features; Beauchemin et al. [26] concluded that a smaller diffusion coefficient provided a better match for the autocorrelation seen in experiments. Sego et al. argued that, compared to diffusion in water, virions were likely to encounter significant additional drag in mucus and therefore suggested that the diffusion coefficient could be as much as six orders of magnitude smaller [183]. Alternatively, You and Yin [219] and Yin and McCaskill [217] have suggested that diffusion is limited by adsorption to cells as virions spread in space. There is very little agreement among the literature as to what the true value of the diffusion coefficient should be under different conditions, yet this value is crucial. Holder et al. [95] and Beauchemin [25] have both demonstrated that the dynamics of spatially–structured viral infections hinge on the value of the viral diffusion coefficient. Williams et al. showed that, in a multicellular model, the relative contributions of cell–to–cell and cell–free infection could only be distinguished for sufficiently large extracellular viral diffusion [209]. Clearly, there is a need to quantify viral diffusion coefficients with increased accuracy. Fortunately, recent advances in imaging technology and analysis have hinted that more accurate quantification might soon be possible. Imle et al. used image tracking to empirically quantify the mean squared displacement and thus diffusion coefficient of HIV virions in different media [99]. The authors determined that the diffusion coefficient in suspension provided reasonable agreement with the Stokes–Einstein estimate, but spread was notably hampered for virions in a collagen matrix [99].
8 Conclusion
This review has demonstrated the broad suite of spatially–structured models of viral dynamics available, and discussed the important contributions they have made to our understanding of viral infections. Advancements in the field point to a important role for spatial structure in viral dynamics, and have offered insights into key mechanisms of virus–host interactions in a spatial context, across a wide range of virus and host types. Nonetheless, there are improvements still to be made. Key questions remain as to how the spatial spread of infections should be modelled in different contexts, and how those models should be implemented. This review highlights work still to be done in better coupling modelling work with experimental data, and in constructing models with application–appropriate implementation. Progress in this regard will result in robust, biologically realistic models, capable of offering important insights into viral illnesses and their treatment.

### v2

@@
 Abstract
-There is growing recognition in both the experimental and modelling literature of the importance of spatial structure to the dynamics of viral infections in tissues.
+There is growing recognition in both the experimental and modelling literature of the importance of spatial structure to the dynamics of viral infections within the host.
 Aided by the evolution of computing power and motivated by recent biological insights, there has been an explosion of new, spatially–explicit models for within–host viral dynamics in recent years.
@@
 Mean–field models are more or less united in the common language of ODEs (possibly with some slight modifications to account for, for example, age structure or stochasticity), yet spatially–structured models span a huge variety of methods: from PDEs to compartmental approaches, to agent–based approaches, or combinations thereof.
-To add to the complexity, the actual nature of the spatial structure of infections by different viruses in different host tissues is also widely varied, depending on whether the infection is in vivo or in vitro, whether the host cells are motile or packed in tissue, how permissive the tissue is to virion diffusion, and so on.
+To add to the complexity, the actual nature of the spatial structure of infections by different viruses in different host tissues is also widely varied, depending on whether the infection is in vivo or in vitro, whether the host cells are motile or packed in tissue, how permissive the extracellular environment is to virion diffusion, and so on.
 There is therefore a need for a wide–reaching, intersectional analysis of the full breadth of spatially–structured viral dynamics models to make sense of the breadth of techniques used and the requirements of different applications.
@@
 Here, we employ a scoping approach to the literature on spatially–structured models of viral dynamics in order to systematically and exhaustively review the full range of approaches used.
+1.1 How to read this paper
 The remainder of this paper is structured as follows.
-In Section 2 we present the methodology used for our scoping review, including inclusion / exclusion criteria.
+In Section 2 we present the methodology used for our scoping review, including the inclusion and exclusion criteria.
 We then present an overview of the key publications in Section 3, before moving on to our detailed review of the field.
-In Sections 4 and 5 we present the state of the field by methodology and virus species respectively, before discussing innately spatially dynamic models in Section 6.
-Finally, in Section 7, we present emerging trends in the future developments of spatial modeling of viral dynamics, before closing the paper in Section 8.
+We do so by first examining the literature from the perspective of the methods and techniques used (in Section 4), and then from the perspective of the biological insights these models have generated (in Section 5).
+We discuss distinct topics in these two sections, and in instances where a publication is cited more than once, it is to discuss different aspects of that publication: for example, its mathematical implementation in one section, and its contribution to biological knowledge in another.
+These sections are structured by model type and viral species modelled, respectively, for easy reference.
+Next, in Section 6, we identify some of the main aspects of viral infection dynamics which have been identified in the literature as being inherently spatially structured, and discuss progress made in the modelling literature in describing and understanding these processes.
+Finally, in Section 7, we present emerging trends for future development of spatially–structured models of viral dynamics, before closing the paper in Section 8.
 2 Methods
@@
 Our criterion for inclusion was studies that introduce or apply a spatially–explicit mathematical or computational model of within–host viral dynamics.
+While many models we found considered the viral infection of tissues within a single host, we also found a number of studies of interest which considered only in vitro systems, or bacterial hosts.
+We deemed these studies relevant and useful to our discussion, and therefore adopted a broad interpretation of “within–host viral dynamics” as the dynamics of viral infection at a between–cell scale.
+Given that neither “within–host” nor “between–cell” featured in our search string, this interpretation will not affect the studies included in the initial search.
 We excluded articles that met one or more of the following criteria:
@@
 We only applied the exclusion criteria at the full text level.
-Of the 3492 unique studies returned by our search, we initially found 92 that were relevant to this review.
+Of the 3939 unique studies returned by our search, we initially found 93 that were relevant to this review.
 The largest subset of publications which were removed on the basis of the exclusion criteria above were articles which were deemed to have an analytical focus (we found 111).
 After having applied our selection criteria, we then screened the references of each of the accepted articles to find any articles missed by our database search.
-This process found a further 84 relevant studies for a total of 176.
+This process found a further 84 relevant studies for a total of 177.
 We outline the general workflow of the literature search process in Figure 1.
@@
 We included pre–prints in our search.
-Figure 1: Workflow for the scoping review. Duplicates automatically detected in JabRef. Duplicates manually detected.
+Figure 1: Workflow for the scoping review.
+We queried Scopus and Web of Science using our search string to find an initial list of potentially relevant studies, then sequentially refined the list of studies based on our inclusion and exclusion criteria.
+We then repeated this approach on the reference lists of the included studies to end up with our final list of accepted publications.
+Duplicates automatically detected in JabRef.
+Duplicates manually detected.
 2.3 Categorisation
@@
 Note that we consider models with a global virus reservoir which has equal access to all spatial regions as having a secondary spatial scale and therefore multicellular.
-We summarise this classification in Figure 2.
-We assign to each study a single modelling framework: for those which make use of multiple models of different frameworks, we categorise the article by the model with the highest level of spatial resolution used (according to our classification, which corresponds to the furthest to the right on the tree in Figure 2).
+We summarise this classification in Figure 2, and provide illustrations of how spatial structure is typically — but by no means exclusively — represented within each model framework.
+We assign to each study a single modelling framework: for those which make use of multiple models of different frameworks, we categorise the article by the model with the highest level of spatial resolution used (according to our classification, which corresponds to the bottom of the tree in Figure 2).
 For each study, we also recorded its number of citations according to Google Scholar as of September 2nd, 2024.
-Figure 2: Classification for modelling frameworks discussed in this review. Note, the boundaries of the methodology classifications used here are maintained through out the reminder of this paper.
+Figure 2: Classification for modelling frameworks discussed in this review.
+For each of the model frameworks discussed in this review (continuum models, patch models, cellular automata and multicellular models) we show an indicative diagram of how space is represented in the system.
+In continuum models, all quantities are modelled as densities in continuous space; in patch models, space is discretised into regions within which cells and other system quantities are represented as densities; in cellular automata, space is discretised into (typically smaller) regions, within which cells are represented by discrete counts.
+Commonly, as we depict here, discrete regions correspond to the cells themselves.
+Multicellular models, like cellular automata, include discretisation of space and track discrete counts of cells (again commonly discretising space into the regions occupied by the cells themselves) but also describe some other spatial scale, such as a continuous virus density.
 3 An overview of the use of spatially–structured models in viral dynamics
-Our search procedure returned 176 publications, spanning 60 years of research across a diverse group of virus species and a wide array of methodologies.
-The final annotated list of publications retained for analysis can be found on FigShare [^1] https://doi.org/10.26188/27085984.v1.
+Our search procedure returned 177 publications, spanning 60 years of research across a diverse group of virus species and a wide array of methodologies.
+The final annotated list of publications retained for analysis can be found on FigShare [^1] https://doi.org/10.26188/27085984.v2.
 In Figure 3 we present a broad overview of the distribution of the studies found by our search over time, and by virus modelled and method used.
@@
 Since 2020 and the outbreak of the COVID–19 pandemic, there has been a spike in the development of spatial models for viral infections, with the average yearly number of publications roughly doubling from less than six in the mid–’10s to approximately 13 in 2023.
-Curiously, this cannot be entirely attributed to models for SARS–CoV–2, with models for the virus accounting for less than a third of recent publications (19 of 62 publications since 2020).
-(a) Refer to caption
-(b) Refer to caption
-(c) Refer to caption
-Figure 3: Spatially–structured viral dynamics models over time (a) by virus species and (b) by methodology. (c) Overall split of models by methodology.
-As shown in Figure 3(c), continuum models formed the largest proportion of the studies found by our search with 56, although between cellular automaton (54) and multicellular (52) approaches, we observed that agent–based models in general were a more popular approach (106 publications combined).
+Curiously, this cannot be entirely attributed to models for SARS–CoV–2, with models for the virus accounting for less than a third of recent publications (19 of 63 publications since 2020).
+Figure 3: Spatially–structured viral dynamics models over time (a) by virus species and (b) by methodology.
+Here we plot the five–year moving average number of publications per year (for 2022 and 2023, we show the average number of publications from two years prior, up to 2023; we do not show data for 2024).
+(c) Overall split of models by methodology.
+The number of publications for each method is given in brackets.
+As shown in Figure 3(c), continuum models formed the largest proportion of the studies found by our search with 56, although between cellular automaton (54) and multicellular (53) approaches, we observed that agent–based models in general were a more popular approach (107 publications combined).
 Early agent–based models were almost exclusively cellular automata, although this approach fell out of favour in the 2010s, where multicellular approaches rose to prominence, as can be seen in Figure 3(b).
@@
 We sought to identify which of the publications found had been the most impactful.
-To account for the time since publication, we quantified impact by considering both a study’s overall number of citations, and the average number of citations it had received per year since publication, and imposed thresholds on both quantities.
-These constraints ensure that a publication has both reached a specified degree of influence overall, but also that it has been influential relative to the number of years since its publication.
+Though clearly not a complete picture of a study’s impact within the field, we chose to quantify this using citation data in order to limit editorial bias in the selection.
+To account for the time since publication, we considered both a study’s overall number of citations, and the average number of citations it had received per year since publication, and imposed thresholds on both quantities.
+These constraints were intended to identify publications that have reached a specified degree of influence overall, but have also been influential relative to the number of years since publication.
 In Figure 4, we plot each publication found in our search against its overall number of publications and its average number of publications per year.
@@
 We opted for thresholds on the overall and per–year citation counts of 50 and 6, respectively, and found 25 studies that passed both thresholds.
-We show a timeline of these seminal publications in Figure 5.
-In the proceeding analysis, we make particular mention of these high–impact publications.
-In this review, articles that were included in our search are referenced alongside the authors’ names, and high–impact publications are indicated in bold.
-Figure 4: Selection of seminal papers.
+We show a timeline of these highly–cited publications in Figure 5.
+In the proceeding analysis, we make particular mention of these highlighted publications.
+In this review, articles that were included in our search are referenced alongside the authors’ names, and highly–cited publications are indicated in bold.
+Figure 4: Identification of papers as highly–cited.
 We show each accepted study for our scoping review by total number of citations and average citations per year since publication.
-Cutoffs for inclusion in our analysis are shown.
-Figure 5: Seminal spatially–structured viral dynamics models over time.
+Cutoffs for definition as highly–cited in our analysis are shown.
+Citation data was collected from Google Scholar on September 2nd, 2024.
+Figure 5: Highly–cited spatially–structured viral dynamics models over time.
 Criteria for inclusion are more than 50 total citations and more than 6 citations per year on average since publication.
 Pre–print.
-Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Wu et al.
+Highly–cited publications are as follows: Bernaschi and Castiglione [24], Zorzenon dos Santos and Coutinho [25], Wu et al.
 [26], Wein et al.
@@
 Figure 6: Continuum models summary.
-Seminal publications are as follows: Wu et al.
+Left: Average number of publications using a continuum model per year.
+Centre: Continuum models by viral species studied.
+Right: Highly–cited continuum model publications over time.
+Highly–cited publications are as follows: Wu et al.
 [26], Wein et al.
@@
 4.2 Patch models
-Figure 7: Patch models summary. Seminal publications are as follows: Funk et al. [75], and Schiffer et al. [177].
+Figure 7: Patch models summary.
+Left: Average number of publications using a patch model per year.
+Centre: Patch models by viral species studied.
+Right: Highly–cited patch model publications over time.
+Highly–cited publications are as follows: Funk et al.
+[28], and Schiffer et al.
+[37].
 Patch models divide the spatial domain into discrete, interconnected regions (or “patches”), within each of which conditions are assumed to be well–mixed.
@@
 Their model considered distinct lymphoid follicles, interconnected via a free blood and lymph compartment, where the spherical lymphoid follicles were in turn subdivided into radial shell patches [95, 93, 94].
+Recent experimental and modelling results have lent credibility to this “patch” representation of secondary lymphoid tissues by identifying clear distinctions in HIV viral load and turnover in follicular and extrafollicular compartments, which may also influence viral evolutionary dynamics [96, 97].
 The within–patch dynamics of patch models need not be governed by deterministic systems.
@@
 Figure 8: Cellular automata summary.
-Seminal publications are as follows: Bernaschi and Castiglione [32], Zorzenon dos Santos and Coutinho [224], Beauchemin et al.
+Left: Average number of publications using cellular automata per year.
+Centre: Cellular automaton publications by viral species studied.
+Right: Highly–cited cellular automaton publications over time.
+Highly–cited publications are as follows: Bernaschi and Castiglione [24], Zorzenon dos Santos and Coutinho [25], Beauchemin et al.
 [29], Xiao et al.
@@
 Figure 9: Multicellular models summary.
-Seminal publications are as follows: Mitchell et al.
+Left: Average number of publications using multicellular models per year.
+Centre: Multicellular models by viral species studied.
+Right: Highly–cited multicellular model publications over time.
+Highly–cited publications are as follows: Mitchell et al.
 [34], Yakimovich et al.
@@
 A similar idea was applied by Howat et al., who considered the spread of free virus at the cell scale, but type I interferon (comprised of molecules far smaller than virions which therefore diffuse much more quickly) as a global quantity [155].
+Similarly, Durso-Cain et al.
+considered a diffusing viral population but a global antibody population [156].
 Models which consider only infection from a global viral reservoir do not give rise to spatially–structured infections, and as such, the global viral population is incorporated into the model dynamics in different ways.
@@
 5.1 HIV
-Figure 10: Spatially–structured models for HIV. Seminal publications are as follows: Zorzenon dos Santos and Coutinho [224], and Imle et al. [99].
+Figure 10: Spatially–structured models for HIV.
+Left: Average number of models for HIV per year.
+Centre: HIV models by model type.
+Right: Highly–cited publications for HIV models over time.
+Highly–cited publications are as follows: Zorzenon dos Santos and Coutinho [25], and Imle et al.
+[2].
 HIV is a chronic infection of lymphoid tissue, particularly CD4+ T cells [172, 92].
@@
 Figure 11: Spatially–structured models for oncolytic viruses.
-Seminal publications are as follows: Wu et al.
+Left: Average number of models for oncolytic viruses per year.
+Centre: Oncolytic virus models by model type.
+Highly–cited publications for oncolytic virus models over time.
+Highly–cited publications are as follows: Wu et al.
 [26], Wein et al.
@@
 This approach suggested that, in order for virotherapy to be effective, the number of viral foci would have to be extremely high, such that the overall infected volume would cover a sufficiently wide region [67].
+However, since the model is phenomenological, the authors cannot deduce a mechanistic explanation for the failure of the treatment at a reasonable dosage.
 This spatially constrained spread of infection foci within tumours is a recurring theme in explaining the limited in vivo efficacy of oncolytic viruses.
@@
 showed that oncolytic infections spread more slowly in a three–dimensional tumour model compared to an equivalent model in two dimensions, despite cells in the former having a higher number of adjacent tumour cells [188].
-Since the spread of viral infections in tumours is mainly between neighbouring cells, the arrangement of cells in space is highly important in agent–based modelling of oncolytic viruses.
+The lack of consensus in these findings points to complex, multifactorial dynamics at play, and suggests a need for an improved understanding of the various mechanisms of oncolytic virus dispersal within tumours, including through cell–to–cell transmission, diffusion, and haptotaxis.
+Moreover, these results suggest that a realistic representation of the tumour microenvironment, such as the extracellular matrix density, the distribution of cancer cells and non–cancer cells, and perhaps the dimensionality of the tumour, may be necessary to ensure dynamics faithful to oncolytic viral infection in vivo.
+Since the spread of viral infections in tumours is believed to mainly between neighbouring cells, the arrangement of cells in space is likely highly important in agent–based modelling of oncolytic viruses.
 Simpler approaches to the arrangement of tumour cells in space consider a regular lattice of cell–sites which remain fixed over time (Wodarz et al.
@@
 compared oncolytic viral dynamics on a regular and Voronoi cell lattice in three dimensions and found that infections on the Voronoi lattice were less likely to result in eradication of the tumour [188].
-Temporal evolution of the cell grid has also been implemented.
-Jenner et al.
+Spatiotemporal evolution of the cell grid is also likely important in the dynamics of oncolytic viruses (as compared to other viral species) due to the rapid growth of cancer tissue.
+Indeed, as we have seen, several authors describing oncolytic viral dynamics at the whole-tumour scale have shown that the growth of the tumour is a significant obstacle to effective treatment (Wein et al.
+[27], Al-Johani et al.
+[184], Simbabwa et al.
+[185], for example).
+However, only one model at the cell scale has accounted for growth of the tissue: Jenner et al.
 considered a time–varying Voronoi lattice model for tumour cells, which also enabled the explicit representation of tumour cell proliferation and remodelling, and moreover facilitated the simulation of time–varying adjacency between cells [165].
+Further work is needed, however, to more deeply explore the dynamics of infection on not just a realistic static cell packing, but also an evolving one.
 5.3 Respiratory viruses
 Figure 12: Spatially–structured models for influenza and SARS–CoV–2.
-Seminal publications are as follows: Beauchemin et al.
+Left: Average number of models for respiratory viruses per year.
+Centre: Respiratory virus models by model type.
+Right: Highly–cited publications for respiratory virus models over time.
+Highly–cited publications are as follows: Beauchemin et al.
 [29], Mitchell et al.
@@
 We note that, to the level of detail that mathematical models have described them, there is little difference between the dynamics of different respiratory viral species; as such in the following we discuss models for respiratory viral infections in general.
-Viruses which infect the respiratory tract specifically target the monolayer of epithelial cells which line the airways [53, 149, 76].
+Viruses which infect the respiratory tract usually target the monolayer of epithelial cells which line the airways [159, 119, 15].
 Thus, to an extent, the natural geometry of respiratory infection is two–dimensional, especially at the microscopic scale, and as a consequence virtually all of the respiratory infection models retained for analysis from our search also consider two spatial dimensions.
-However, at a larger length scale, there is substantial heterogeneity in the structure of the lung along the length of the respiratory tract.
+However, at a larger length scale, there is substantial structural heterogeneity along the length of the respiratory tract.
 The respiratory tract has a complex, tree–like structure, comprising 24 generations of branching airways, such that upper passages like the trachea are relatively wide (on the order of centimetres), while deeper airways are very narrow (on the order of fractions of millimetres) [189, 159].
@@
 While the use of this hyperrealistic geometry demonstrates the notable differences in the dynamics of infections on flat sheets compared to the lung, there is a substantial conceptual leap between the two geometries.
-Other authors have instead employed an intermediate approach, studying the spread of viral infections on sections of the lung in order to identify specific ways in which the structure and geometry of lung tissue influence the spread of infection.
+Other authors have instead employed an intermediate approach, studying the spread of viral infections on sections of the respiratory tract in order to identify specific ways in which the structure and geometry of the respiratory tract influence the spread of infection.
 Chen et al.
 [159] and Williams et al.
-[208] studied the spread of viral infections on individual lung branches, or on small sections of branching tissue.
+[191] studied the spread of viral infections on individual airway branches, or on small sections of branching tissue.
 The latter group showed that tubular and branching tissue geometry imposes rich, spatially–directional infection dynamics [191].
 These dynamics moreover exhibited an increased sensitivity to the rate of viral diffusion compared to the dynamics of infection on a flat cell sheet [191].
+While these theoretical works have offered hints into the ways in which the dynamics of respiratory viral infections in vivo may differ from infections in vitro, a deeper integration with experimental data is now necessary to gain quantitative insights into the spread of infection in the respiratory tract.
+Obtaining relevant in vivo data remains a challenge, however, advances in imaging techniques and genetic engineering (discussed in Section 7.2) may soon provide data necessary for such analysis.
 In addition to structural heterogeneity, there is also variability in the cellular composition of upper and lower regions of the respiratory tract.
@@
 [43] included the mucociliary escalator in a model of influenza infection, and suggested that infection can effectively only spread up the respiratory tract from the initial point of virion deposition [43].
-The rate of mucociliary flow is variable at different depths of the respiratory tract, with slower flow in deeper regions of the lung [130, 53].
+The rate of mucociliary flow is variable at different depths of the respiratory tract, with slower flow in lower airways [189, 159].
 Chen et al.
@@
 Figure 13: Spatially–structured models for other specific viruses.
-Seminal publications are as follows: Xiao et al.
+Left: Average number of models for other specific viruses per year.
+Centre: Other specific virus models by model type.
+Right: Highly–cited publications for other virus models over time.
+Highly–cited publications are as follows: Xiao et al.
 [31], Heilmann et al.
@@
 Bacteria and phage inhabit and interact within a broad array of environments, including in soil, water, and within larger organisms [33, 152].
-We note that there is a major difference between phage dynamics and the dynamics of the other virus species which we have considered in this review, in that phage hosts are single–cell organisms.
-As such, while models for phage–bacteria interactions describe the spread of infection among cells — as with the other models we have discussed — such systems are, strictly, between–host models.
-We have nonetheless included such studies in our analysis.
-Due to the unique properties of phage–bacteria dynamics, as compared to the dynamics of other viral species we have discussed, modellers have considered the spatially–structured interactions between phage and microbe communities on far larger spatiotemporal scales.
+There is a major difference between phage dynamics and the dynamics of the other virus species which we have considered in this review, in that while we have so far mainly considered viruses which infect tissues within a single host, the bacteria which phage infect are single–cell organisms.
+However, there are also many similarities.
+Phage–bacteria systems are frequently studied — and modelled — in vitro (Valdez et al.
+[194], de Rioja et al.
+[50], for example), and the dynamics of these systems can be directly compared to the in vitro dynamics of other, human–adapted viruses.
+Bacteria also frequently form dense, surface–bound microbial communities called biofilms, which can be likened to the crowded environment within tumours [194, 44, 41].
+Despite first appearances, modelling of phage–bacteria dynamics is therefore closely interlinked with models for other viral species, and forms an important component of the development of spatially–structured viral dynamics models in general.
+Due to the unique properties of phage–bacteria dynamics, as compared to the dynamics of other viral species we have discussed, modellers have considered the spatially–structured interactions between phage and microbe communities on very large spatiotemporal scales.
 Garrido Zornoza et al.
@@
 The authors showed that this influx–outflux process among habitats lead to the spatial spread of microbial communities but also resulted in destabilisation and occasional local extinctions in individual habitats [152].
-While phage–bacteria systems exhibit spatially–structured dynamics on this ecological spatial scale, most authors instead model the dynamics of phage–bacteria interactions on far smaller length scales.
-In particular, many authors have modelled the spread of phage infection among dense, surface–bound microbial communities called biofilms.
+While phage–bacteria systems exhibit spatially–structured dynamics on this ecological spatial scale, most authors instead model the dynamics of phage–bacteria interactions on far smaller length scales, particularly in biofilms.
 The crowded nature of these environments requires particular consideration by modellers.
@@
 Taylor et al., for example, showed that the high cell density found in biofilms promotes the formation of spatial clustering of infection [196].
+Valdez et al.
+showed that, as a consequence of this highly localised spread, differently–shaped initial distributions of phage among bacterial communities gave rise to spatially distinct colony structures [194].
 Modellers have accounted for the crowded nature of biofilms in various ways.
-While some, including Heilmann et al.
+Some, including Heilmann et al.
 (2012) [36] and Mitarai et al.
@@
 explicitly considered physical blocking of phage motion by bacteria [44].
-Valdez et al.
-showed that, as a consequence of this highly localised spread, differently–shaped initial distributions of phage among bacterial communities gave rise to spatially distinct colony structures [202].
+A direct comparison of these different potential mechanisms is lacking, but could be compared to experimental evidence to provide insight into the characteristics of phage dispersal in the crowded biofilm environment, which may also have consequences for other viral species, especially oncolytic viruses.
 One of the remarkable features of phage–bacteria interactions is the coexistence of both populations over vast time scales, even within a single biofilm.
@@
 showed that viral cccDNA represents one of the most robust obstacles to hepatitis B treatment, and suggested that even a single cccDNA molecule would be sufficient to reinfect the entire liver [153].
+Interestingly, although the non–spatial modelling literature for hepatitis B and C is fairly extensive (reviewed by Dahari et al.
+[202] and Canini and Perelson [203], for example), there are relatively few (seven) articles which account for spatial structure.
+This is despite clear biological evidence of the role of spatial dynamics in the spread of hepatitis B and C infection, such as the presence and suspected high rate of cell–to–cell transmission [204, 99, 107], and experimentally observed clusters of infected cells in hepatitis C–infected livers [203, 205].
+There is therefore substantial scope for future model development for hepatitis viruses which takes into account spatial effects in the dynamics.
 Some authors have also modelled the spatial dynamics of herpes simplex virus (HSV–2).
@@
 In this section, we review how spatially–structured models have accounted for spatially–acting aspects of viral infection, and discuss how these dynamics differ from descriptions by non–spatial models.
+Figure 14 illustrates the sources of spatially–structured viral dynamics discussed in this section in a biological context.
+Figure 14: Sources of spatially-structured viral dynamics discussed in Section 6.
 6.1 Spatial heterogeneity and parameter estimation
 One of the simplest improvements offered by spatially–structured modelling is in allowing for spatially–heterogeneous distribution of infection.
+Heterogeneity of the host tissue is illustrated in Figure 14.
 We have already discussed several examples of heterogeneous tissues, including heterogeneity in pressure and density within tumours for oncolytic viruses (Morselli et al.
@@
 However, there has been growing recognition, both in the biological and modelling literature, of the importance of the cell–to–cell route of infection [209, 210, 211, 212, 15].
+Cell–to–cell transmission is illustrated in the bottom left of Figure 14.
 This mode of transmission has been observed in infections by virtually all medically–important viruses, including retroviruses, herpesviruses, coronaviruses and influenza viruses [212, 204].
@@
 fit a multicellular model to experimental data for hepatitis C and concluded that cell–to–cell infection could account for as much as $99$% of infection events, however cell–free infection was crucial to the overall progression of the virus by providing access to uninfected regions of the tissue away from the infection focus [107].
-Goyal and Murray, by contrast, argued that cell–to–cell infection was not especially important to the overall dynamics of hepatitis B in the early stages of infection, but was crucial in the chronic stage of infection through T cell exhaustion [82].
+Durso-Cain et al., who also modelled hepatitis C dynamics, reported a similar synergistic effect between the two modes of transmission, and also noted that the cell–to–cell mechanism was dominant, especially early in the infection [156].
+By contrast, Goyal and Murray, argued that in hepatitis B, cell–to–cell infection was not especially important to the overall dynamics in the early stages of infection, but was crucial in the chronic stage of infection through T cell exhaustion [99].
 On this longer time scale, Saeki and Sasaki [3] and Kreger et al.
@@
 In addition to the spatial spread of infection, there are also aspects of the immune response which are inherently spatially–structured.
+Immune system dynamics in the context of viral infection are illustrated in the upper portion of Figure 14.
 Modelling has shown that the presence of spatial structure significantly increases the potency of the interferon–mediated innate immune response (Lavigne et al.
@@
 Here, we have conducted a scoping review to show the huge diversity in approaches to modelling spatial aspects of viral infections.
-The wide range of methods and approaches used mirrors the significant diversity in modelling requirements between applications, including particularities of viral species or model tissues, for example.
+The wide range of methods and approaches used mirrors the significant diversity in modelling requirements between applications, including particularities of viral species or host cell populations, for example.
 We have shown that spatial models offer a description of a range of biological phenomena with an intrinsically spatial mode of action, such as cell–to–cell transmission, where an accurate description is not possible using traditional ODE modelling.
@@
 [29] have both shown that stochastic properties of cellular automaton models are influenced by the size of the grid.
+In another study, Williams et al.
+studied multicellular models in which the host tissue is discretised into cells and the virus population is represented by a density surface, and showed that for rates of viral diffusion akin to standard assumptions in the literature, the dynamics of the system was dependent on the resolution of spatial discretisation of the viral surface [171].
 Berg et al.
@@
 [119] have also demonstrated that the dynamics of their models varied significantly depending on whether two or three spatial dimensions were considered.
-This finding is not necessarily a limitation of the agent–based approach: the dependence of model dynamics on dimensionality, for example, is an important biological finding in its own right.
-Nonetheless, this implementation–dependence clearly highlights the importance of biological fidelity in model construction, and moreover points to a need for careful study of various elements of the construction of agent–based models and how they might influence the dynamics.
-Key questions, such as the spatial arrangement of cells, the size and dimensionality of the spatial domain, or the discretisation of computational multicellular systems, have so far been largely unexplored in viral dynamical models, yet are crucial for ensuring robust implementation of agent–based systems.
+Not all of these findings are necessarily limitations of the agent–based model framework: the dependence of model dynamics on dimensionality, for example, is an important biological finding in its own right.
+This implementation–dependence clearly highlights the importance of biological fidelity in model construction, and moreover points to a need for careful study of various elements of the construction of agent–based models and how they might influence the dynamics.
+Key questions, such as the spatial arrangement of cells, the size and dimensionality of the spatial domain, or the numerical simulation of complex, multi–scale systems, have so far been largely unexplored in viral dynamical models, yet are crucial for ensuring robust implementation of agent–based models.
+Another emerging trend in spatially–structured modelling is the growing importance of computational implementation.
+Recent model development, especially of large, complex systems, has increasingly shifted modelling language away from the mathematical and towards a software–oriented approach.
+In particular, we noted in Section 4.3 the increasing popularity of large viral dynamics “simulators”, the development of which is usually facilitated by existing software packages for tissue dynamics such as Chaste [130, 129], PhysiCell [123], Morpheus [131] and others.
+This direction in model development, on the one hand, is an important step towards modularity and submodel integration (illustrated especially in the work of Getz et al.
+[9, 120]), but also poses challenges in model interpretability and reproducibility, including between software platforms.
+One aspect of this is because different software packages incorporate different dynamics which are “baked in” to the base tissue dynamics model, such as the Cellular Potts Model in CompuCell3D [8, 125, 124, 126, 127] or the point force cell mechanics model of PhysiCell [9, 120, 123].
+This complicates comparison of models between software platforms, even if the assumed virus–host interactions are the same.
+In other fields of mathematical biology (systems biology in particular), efforts have increasingly been made to standardise and formalise modelling language through initiatives like Systems Biology Markup Language (SBML) [225], CellML [226] and BioNetGen [227].
+While none of these approaches apply easily to spatially–structured systems (especially agent–based systems), as models grow increasingly complex and software–dependent, future development in standardisation approaches would aid in the communication and reproducibility of spatially–structured models of viral dynamics and of biological systems more generally.
 7.2 Application to experimental data
 Perhaps the greatest obstacle to the wider usage of spatially–structured models of viral dynamics is the difficulty in calibrating such models to experimental data.
-Spatial models are inherently more complex and computationally expensive than their non–spatial counterparts and therefore both obtaining sufficient spatial data and carrying out rigorous parameter estimation presents something of a challenge [85, 86, 24, 76].
-While some authors have conducted such a fitting process using spatial models (Holder et al.
-[96], Haseltine et al.
-[92]), the vast majority instead rely on drawing at least some parameter estimates from elsewhere in the literature.
+Spatial models are inherently more complex and computationally expensive than their non–spatial counterparts and therefore both obtaining sufficient data as well as carrying out rigorous parameter estimation presents a significant challenge [11, 13, 10, 15].
+Perhaps as a consequence, in the vast majority of spatially–structured models, authors rely on drawing at least some parameter estimates from elsewhere in the literature.
 However, the exact meaning of a particular parameter is likely to vary somewhat between models depending on what components and interactions are included.
@@
 The authors determined that the diffusion coefficient in suspension provided reasonable agreement with the Stokes–Einstein estimate, but spread was notably hampered for virions in a collagen matrix [2].
+At least qualitatively, this confirms the assumption of slowed viral diffusion in vivo, but moreover hints at the limited applicability of in vitro viral spread in a within–host context.
+The mismatch between in vitro observations and viral dynamics in vivo is a broader challenge in the development of biologically relevant spatially–structured models.
+Data describing the spatial spread of viral plaques in vitro has been readily available for decades [22], and has been successfully used to fit a number of spatially–structured models (Holder et al.
+[53], Haseltine et al.
+[56], Imle et al.
+[2], for example).
+However, as we mentioned above, the lack of realistic viral diffusion in the in vitro context, as well as a simplified tissue structure and the lack of an immune response, means that such models may be of limited relevance to in vivo dynamics.
+While spatially–structured data from infections in vivo remains scarce, emerging genetic engineering and imaging technologies suggest possible new sources of information about how infections spread within living hosts.
+Several groups have engineered reporter viruses expressing fluorescent proteins, which can be detected by bioluminescence imaging in animal hosts, permitting the construction of precise maps of viral distribution in vivo at various times over the course of infection [232, 233, 234].
+Other groups have developed reporter mice which express fluorescent proteins as part of, for example, the innate immune response [235, 236].
+This likewise permits imaging of interferon distribution, for example, during viral infection.
+Other techniques, such as FDG-PET imaging, permit the visualisation of the extent of infection within living hosts (including human hosts) without the need for an engineered reporter virus, albeit at a lower resolution [237, 78].
+While these technologies provide useful descriptions of how the infection distribution evolves within the host, currently, data is usually available only at the macroscopic scale.
+In limited cases, some spatially–structured data can be gathered at the cell scale, but requires removing samples of infected tissue, and thus sacrificing the host [77, 76].
+As a consequence, it is not possible to collect multiple such measurements over the time course of a single infection.
+Nonetheless, such data may still be valuable in calibrating realistic spatially–structured models of viral infections in vivo.
+One promising publication by Fukuyama and colleagues studied mice co-infected with four engineered influenza strains, each expressing differently–coloured fluorescent proteins, and examined samples of lung tissue taken at different stages in the infection [76].
+Imaging showed not only temporal progression of the infection from the upper to the lower airways, but showed local clustering of infected cells with the same fluorescent tag, suggesting competition between local and distal infection mechanisms [76].
+Emerging techniques such as these can therefore offer rich sources of spatiotemporal information about the spread of viral infection within the host, which will further aid in the development and calibration of biologically realistic models.
+A more practical obstacle in the calibration of spatially–structured models to data is in ensuring computational tractability of parameter fitting methods.
+Spatially–structured models are typically computationally expensive to simulate, which necessitates sophisticated, highly efficient parameter estimation methods — often coupled with high–performance computing — in order to conduct model calibration in feasible times [10, 238].
+Computational demands are even greater for stochastic models, for which it may be more difficult to determine the quality of fit for a given set of parameter values [239, 240].
+There are now a range of inference methods and software tools available for these parameter estimation tasks, including Bayesian and Approximate Bayesian Computation (ABC) methods which quantify parameter uncertainty (Stan [241], pyABC [242], ABCpy [243], for example).
+However, a barrier to wider uptake of this approach is a need for specific technical expertise in implementing and interpreting such methods, as well as a lack of adaptability in existing implementations [238].
+However, just as there have been efforts to unify modelling languages under standards like SBML, several authors have also recently sought to introduce standards for parameter inference problems.
+PEtab, in particular, is a popular such format for parameter estimation on SBML models [244].
+Promisingly, efforts have also recently been made to introduce standard workflows for parameter inference on spatially–structured models in particular, such as Spatial Model Editor [245] and FitMultiCell [238].
+Greater accessibility and interpretability of these advanced parameter estimation techniques will lead to better integration of experimental data, even with complex spatial models, as well as aid in reproducibility of results.
 8 Conclusion
@@
 Progress in this regard will result in robust, biologically realistic models, capable of offering important insights into viral illnesses and their treatment.
+Acknowledgements
+TW’s research is supported by an Australian Government Research Training Program (RTP) scholarship.
+JMM’s research is support by the ARC (DP210101920).
+JMO’s research is supported by the ARC (DP230100380, FT230100352).
+The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
+The images in Figure 14 were created in BioRender.
+Osborne, J.
+(2025) https://BioRender.com/t2kwkm3.